### v2

# Volume and topology of bounded and closed hyperbolic $3$-manifolds, II

Abstract.
Let $N$ be a compact, orientable hyperbolic 3-manifold whose boundary is a connected totally geodesic surface of genus $2$. If $N$ has Heegaard genus at least $5$, then its volume is greater than $2V_{\rm oct}$, where $V_{\rm oct}=3.66\ldots$ denotes the volume of a regular ideal hyperbolic octahedron in ${\mathbb{H}}^{3}$. This improves the lower bound given in our earlier paper “Volume and topology of bounded and closed hyperbolic $3$-manifolds.” One ingredient in the improved bound is that in a crucial case, instead of using a single “muffin” in $N$ in the sense of Kojima and Miyamoto, we use two disjoint muffins. By combining the result about manifolds with geodesic boundary with the $\log(2k-1)$ theorem and results due to Agol-Culler-Shalen and Shalen-Wagreich, we show that if $M$ is a closed, orientable hyperbolic $3$-manifold with $\mathop{\rm vol}M\leq V_{\rm oct}/2$, then $\dim H_{1}(M;{\mathbb{F}}_{2})\leq 4$. We also provide new lower bounds for the volumes of closed hyperbolic $3$-manifolds whose cohomology ring over ${\mathbb{F}}_{2}$ satisfies certain restrictions; these improve results that were proved in “Volume and topology….”
1. Introduction
It follows from the Mostow rigidity theorem that the volume of a finite-volume hyperbolic $3$-manifold $M$ is a topological invariant of $M$. We may regard $\mathop{\rm vol}M$ as a measure of the topological complexity of $M$, and it is natural to try to relate this measure of topological complexity with more classical ones, such as the ranks of homology groups of $M$ with prescribed coefficients.
It has long been known that for any prime $p$, the dimension of $H_{1}(M;{\mathbb{F}}_{p})$ (where ${\mathbb{F}}_{p}$ denotes the field of order $p$) is linearly bounded in terms of $\mathop{\rm vol}M$. According to [19, Theorem 5.4], we have $\dim H_{1}(M;{\mathbb{F}}_{p})<168.602\cdot\mathop{\rm vol}M$ for every prime $p$. It is expected that in the forthcoming paper [20] this result will be improved in the case $p=2$, by replacing the coefficient $168.602$ by one that is a bit less than $158$.
For small values of $\mathop{\rm vol}M$, these results were improved by a couple of orders of magnitude in [2], [11], [12], and [18]:
  • Theorem 1.1 of [2] asserts that if $\mathop{\rm vol}M\leq 1.22$ then $\dim H_{1}(M;{\mathbb{F}}_{p})\leq 2$ for $p\neq 2,7$, while $\dim H_{1}(M;{\mathbb{F}}_{p})\leq 3$ if $p$ is $2$ or $7$;
  • Theorem 1.2 of [11] asserts that if $\mathop{\rm vol}M\leq 3.08$ then $\dim H_{1}(M;{\mathbb{F}}_{2})\leq 5$.
  • Theorem 14.5 of [18] (which improves on Theorem 1.7 of [12]) asserts that if $\mathop{\rm vol}M\leq 3.69$ then $\dim H_{1}(M;{\mathbb{F}}_{2})\leq 7$.
Among these results, only Theorem 1.1 of [2] is known to be sharp, and only for $p=5$: we have $\dim H_{1}(M;{\mathbb{F}}_{5})=2$ when $M$ is the Weeks manifold. Among the closed, orientable hyperbolic $3$-manifolds $M$ of volume at most $3.69$ known from a census compiled by Hodgson–Weeks (see [23]) and shipped with SnapPy [9], the largest value of $\dim H_{1}(M;{\mathbb{F}}_{2})$ is $3$; it is achieved by the manifold denoted by ${\rm m135}(-1,3)$ in the census,
In this paper, we will denote by $V_{\rm oct}$ the volume of a regular ideal hyperbolic octahedron. It is known that $V_{\rm oct}=8\El(\pi/4)=3.6638\ldots$, where is the Lobachevsky function; see the Example at the end of Section 7.2 of Thurston’s notes [35]. By the Fourier series for the Lobachevsky function given in [35, Lemma 7.1.2], $V_{\rm oct}$ is equal to four times Catalan’s constant.
We will prove:
Theorem 10.1.
Let $M$ be a closed, orientable hyperbolic $3$-manifold with $\mathop{\rm vol}M\leq V_{\rm oct}/2$. Then $\dim H_{1}(M;{\mathbb{F}}_{2})\leq 4$.
This result improves on Theorem 1.2 of [11] in the range in which it applies, but it is probably not sharp. Among known closed, orientable hyperbolic $3$-manifolds $M$ of volume at most $V_{\rm oct}/2$, the largest value of $\dim H_{1}(M;{\mathbb{F}}_{2})$ is $2$; it is achieved by the manifold denoted by ${\rm m}009(5,1)$ in the census cited above, which is arithmetic and has volume exactly $V_{\rm oct}/2$.
Contrapositively, the results quoted above from [2], [11], and [12] may be interpreted as saying that lower bounds on the dimension of $H_{1}(M;{\mathbb{F}}_{2})$, where $M$ is a closed, orientable, hyperbolic manifold, give lower bounds on $\mathop{\rm vol}M$. In [14] it was shown that these lower bounds become stronger if one places restrictions on the cup product pairing from $H^{1}(M;{\mathbb{F}}_{2})\otimes H^{1}(M;{\mathbb{F}}_{2})$ to $H^{2}(M;{\mathbb{F}}_{2})$. Specifically, Theorem 1.2 of [14] asserts that if $\dim H^{1}(M;{\mathbb{F}}_{2})$ is at least $5$, and the dimension of the cup product pairing is at most $1$, then $\mathop{\rm vol}M>3.44$. In this paper we will prove the following stronger result:
Theorem 11.3.
Let $M$ be a closed, orientable hyperbolic $3$-manifold. Set $r=\dim H_{1}(M;{\mathbb{F}}_{2})$, and let $t$ denote the dimension of the image of the cup product pairing $H^{1}(M;{\mathbb{F}}_{2})\otimes H^{1}(M;{\mathbb{F}}_{2})\to H^{2}(M;{\mathbb{F}}_{2})$. Then:
  1. (1)
    if $r\geq 5$ and $t\leq 1$, we have $\mathop{\rm vol}(M)>3.57$; and
  2. (2)
    if $r\geq 6$ and $t\leq 3$, or $r\geq 7$ and $t\leq 5$, we have $\mathop{\rm vol}(M)>V_{\rm oct}$.
The proofs of Theorems 10.1 and 11.3 use the following result, Proposition 9.4. Recall that a group $\Pi$ is said to be $k$-free for a given positive integer $k$ if every subgroup of $\Pi$ whose rank is at most $k$ is free.
Proposition 9.4.
Let $M$ be a closed, orientable, hyperbolic $3$-manifold, let $k\geq 3$ be an integer, and suppose that $\dim H_{1}(M;{\mathbb{F}}_{2})\geq\max(3k-4,6)$. Then either $\pi_{1}(M)$ is $k$-free, or $\mathop{\rm vol}M>2V_{\rm oct}$.
The proofs of Theorems 10.1 and 11.3 and Proposition 9.4 involve combining a rich variety of topological and geometric techniques. One key ingredient in the proof of Proposition 9.4 is the following result, which pertains to manifolds with boundary and features a different measure of topological complexity, the Heegaard genus, which in this paper will be denoted $\mathrm{Hg}(N)$. For a compact, connected, orientable $3$-manifold NN, with or without boundary, ${\rm Hg}(N)$ is the smallest genus of a closed surface in $N$ that divides it into two compression bodies.
Theorem 7.2.
Let $N$ be a compact, orientable hyperbolic $3$-manifold with $\partial N$ connected, totally geodesic, and of genus $2$. If $\mathrm{Hg}(N)\geq 5$ then $\mathrm{vol}(N)>2V_{\rm oct}$.
This directly strengthens Theorem 1.1 of [14], which gives a lower bound of $6.89$ on the volume of $N$ with the same topological hypothesis. Like the results above for closed manifolds, it is still likely not sharp: in a census compiled by Frigerio-Martelli-Petronio [17], all examples with volume at most $2V_{\rm oct}$ have Heegaard genus $3$. This census collected all manifolds with totally geodesic boundary that decompose into at most four truncated tetrahedra. After Kojima–Miyamoto’s minimal-volume examples (of volume $6.45$...), the census of [17] contains six manifolds with volume $7.10$..., and the next-smallest have volume $7.33...~>2V_{\rm oct}$.
We now describe the structure of the body of the paper, which breaks naturally into three parts. The first of these, consisting of Sections 2 through 5, primarily addresses hyperbolic $3$-manifolds with totally geodesic boundary. The results of these sections build on the “geometric part” of the proof scheme of [14, Th. 1.1], which itself built on vocabulary and results established by Kojima–Miyamoto in their work [27] that identified the minimum-volume compact hyperbolic $3$-manifolds with totally geodesic boundary. Key tools of analysis include return paths and $(i,j,k)$-hexagons, which are respectively associated to pairs and triples of boundary components of the universal cover $\widetilde{N}$ (definitions in Subsection 2.1).
Section 2 reviews the methods introduced in [27] and further developed in [14] for describing how the lengths of return paths of a manifold $N$ satisfying the hypotheses of Theorem 7.2 are controlled by the geometry of $\partial N$, moderated by $(i,j,k)$-hexagons. In places, we incrementally improve these methods. In particular, Proposition 2.2.2 simplifies the lower bound on the second-shortest return path length $\ell_{2}$ as a function of $\ell_{1}$ given in [14, Lemma 2.9] (which itself built on [27, §4]). And Proposition 2.3.1 gives a stronger lower bound than [14, Prop. 3.9] on $\cosh\ell_{1}$ in the absence of a $(1,1,1)$-hexagon: $1.23$ here versus $1.215$ there. (The minimum possible value for $\cosh\ell_{1}$ is $1.183..$, proved in [27].)
Section 3 introduces a significant new tool for bounding volume below: a second muffin. Here, “muffin” is Kojima-Miyamoto’s term for a hyperbolic solid of rotation generated by a certain pentagon with four right angles. They show in [27, Lemma 3.2] that such a muffin embeds in $N$ with its rotation center along the shortest return path $\lambda_{1}$. We called this $\mathrm{Muf}_{\ell_{1}}$ in [14] and continue to do so here. This is the “first” muffin in $N$.
In Subsection 3.1 we define a broader class of muffins and lay out criteria for embedding a second muffin in $N$, centered on the second-shortest return path, without overlapping $\mathrm{Muf}_{\ell_{1}}$. Each muffin that we use intersects $\partial N$ in the union of two disjoint disks, its “caps”. In Subsection 3.2 we give a sufficient condition to ensure that a collar of the region of $\partial N$ outside the muffin caps is embedded in $N$ and does not overlap the muffins.
Section 4 uses embedded muffins and collars to give lower bounds on volume. The fundamental volume bound given in [27] and used in [14], recorded here in (4.2.1), is a function of $x=\cosh\ell_{1}$ that records the sum of the volumes of $\mathrm{Muf}_{\ell_{1}}$ and of an embedded collar of the complement in $\partial N$ of its caps. For manifolds satisfying certain conditions on $\ell_{1}$ and $\ell_{2}$, we bring a second muffin into play here, allowing us to recover additional volume. We prove:
Corollary 4.7.
Let $N$ be an orientable hyperbolic $3$-manifold with $\partial N$ compact, connected, totally geodesic, and of genus $2$. If the universal cover $\widetilde{N}$ of $N$ contains no $(1,1,1)$-hexagon then $\mathrm{vol}(N)\geq 7.4$.
This is the main result of Section 4. It strengthens the lower bound of $6.89$ for $\mathrm{vol}(N)$ which, under the same hypotheses, follows from Propositions 3.7 and 3.9 of [14].
In Section 5 we review and slightly upgrade certain results of [14] that address the other case, in which there is a $(1,1,1)$-hexagon in $\widetilde{N}$. Our upgrades here remove or relax restrictions on $\ell_{1}$ in the hypotheses of their antecedents. Notably, Lemma 5.1 shows unconditionally that $(1,1,1)$-hexagons interact well with shortest return paths, removing a hypothesis of [14, Lemma 6.6]. The thrust of this section follows that of [14, §6], using the $(1,1,1)$-hexagon to construct a submanifold $X$ of $N$ which is a“non-degenerate trimonic” submanifold in the sense defined in [14, §5]). Results from the “topological part” of the proof of [14, Th. 1.1], which are laid out in Sections 4 and 5 of that paper, will then be applied to $X$ without requiring further adaptation.
Sections 6 and 7 constitute the second of the present paper’s three parts. Section 6 first introduces topological notation used in the rest of the paper, in Subsection 6.1. In Subsection 6.2 we refine methods of Agol-Storm-Thurston [4] for bounding the volume of a hyperbolic Haken $3$-manifold $M$ below in terms of the topology of the manifold obtained by cutting $M$ along an incompressible surface. The subsection’s main result, Theorem 6.2.8 strengthens the conclusion of Theorem 9.1 of [4] for compact such $M$ by replacing a non-strict inequality with a strict one; it also allows $M$ to have connected, totally geodesic boundary.
In Section 7 we first prove Theorem 7.1, which improves Theorem 7.4 of [14]. The improvement comes from applying Theorem 6.2.8 (replacing the previous result’s appeal to [4, Th. 9.1]) to the frontier in $N$ of the trimonic submanifold $X$ constructed in Section 5, in a certain case of the proof of this result. The proof of Theorem 7.2 completes the section, with complementary cases supplied there by Theorem 7.1 and Corollary 4.7.
In the paper’s third part, beginning with Section 8, we shift our focus to closed manifolds. Section 8 gives background necessary to prove our results in this setting. The proof of Proposition 9.4 begins by using topological results due to Culler and Shalen [11], about desingularization of $\pi_{1}$-injective singular surfaces in $3$-manifolds, to show that if $M$ satisfies the homological hypothesis of the proposition and $\pi_{1}(M)$ is not $k$-free, then $M$ contains a closed incompressible surface $S_{0}$ of some genus $g$ with $2\leq g\leq k-1$. The homological hypothesis implies that the Heegaard genus of $M$ is strictly greater than $2g+1$.
If $S\subset M$ is any closed incompressible surface, and $M^{\prime}=M\setminus\setminus S$ denotes the manifold obtained from $M$ by splitting it along $S$, we denote by $\mathop{\rm kish}(M^{\prime})$ (sometimes called the “kishkes” or “guts” of $M^{\prime}$) the union of those components of $\overline{M^{\prime}-\Sigma}$ that have strictly negative Euler characteristic, where $\Sigma$ denotes the union of the characteristic submanifolds of the components of $M^{\prime}$.
The existence of an incompressible surface $S_{0}$ of genus $g$, together with the strict lower bound $2g+1$ for the Heegaard genus of $M$, is used—via a result proved in [18] using topological ideas developed by Culler, DeBlois and Shalen in [8]—to produce a closed incompressible surface $S\subset M$ such that either (1) the Euler characteristic $\chi(\mathop{\rm kish}(M\setminus\setminus S))$ is at most 2-2, or (2) $S$ separates $M$, and $M\setminus\setminus S$ has a component which is acylindrical, i.e. contains no essential annulus. If (1) holds, the geometric methods developed by Agol, Storm and Thurston in [4] (via Theorem 6.2.8 of this paper) give a strict lower bound of $2V_{\rm oct}$ for $\mathop{\rm vol}M$.
If (2) holds but (1) does not, and if we fix an acylindrical component $A$ of $M\setminus\setminus S$, then $A$ is diffeomorphic to a hyperbolic $3$-manifold NN with totally geodesic boundary, and the methods of [4] show that $\mathop{\rm vol}N$ is a lower bound for $\mathop{\rm vol}M$. Furthermore, in this case a Mayer-Vietoris calculation shows that $\dim H_{1}(N,{\mathbb{F}}_{2})$, and hence the Heegaard genus of $N$, is at least $5$. If $S$ has genus $2$, Theorem 7.2 now gives a strict lower bound of $2V_{\rm oct}$ for $\mathop{\rm vol}N$. If $S$ has genus greater than $2$, the geometric results established by Miyamoto in [30] give a stronger lower bound.
[…]
The proof of Theorem 11.3 follows the same basic outline as the proof of Theorem 1.2 of [14]. As in the latter proof, we distinguish the cases in which $\pi_{1}(M)$ is or is not $4$-free. If $\pi_{1}(M)$ is $4$-free, one of the main results proved by Guzman and Shalen in [18] gives a lower bound of $3.57$ for $\mathop{\rm vol}M$, which is a surprising improvement over the lower bound of $3.44$ established in [12] and quoted in [14], and is enough to prove Assertion (1) of Theorem 11.3 in this case. If we combine the assumption of $4$-freeness with a lower bound of $6$ for $\dim H_{1}(M;{\mathbb{F}}_{2})$, then arguments given in [18], based on Agol and Dunfield’s results on the change of volume under Dehn drilling [4] and results due to Culler and Shalen about volumes and homology of one-cusped manifolds [10], allow one to obtain a lower bound of $3.69$ for $\mathop{\rm vol}M$, which establishes Assertion (2) of Theorem 11.3 in this case.
[…]
We are grateful to Nathan Dunfield for explaining the example ${\rm m}009(5,1)$ that was referred to above, and to Joel Hass for explaining material related to Proposition 6.2.7 to us.
2. Existing foundations
This section reviews an approach developed by Kojima–Miyamoto [27], and further exploited in our earlier paper [14], to controlling the geometry of a hyperbolic $3$-manifold NN with totally geodesic boundary via the geometry of $\partial N$. We focus on the orthospectrum of $N$, the sequence of lengths of properly immersed arcs in $N$ that meet $\partial N$ perpendicularly.
Section 2.1 relates the orthospectrum of $N$ to a certain spectrum of arc lengths on $\partial N$ using hyperbolic trigonometry. In Section 2.2 we describe Kojima–Miyamoto’s packing arguments for bounding $\ell_{2}$, the second-smallest ortholength, below in terms of $\ell_{1}$. We conclude in Section 2.3 with an observation from [14] that bounds $\ell_{2}$ above in terms of $\ell_{1}$ in the absence of “$(1,1,1)$-hexagons” (see below).
In fact we make incremental improvements below to the existing bounds we describe. These bounds are for the most part still not strong enough to be directly useful to our volume estimates, but we use them to delimit a search space for a procedure to find sharper ones.
2.1. Ortholengths vs arclengths
If $N$ is a complete hyperbolic $3$-manifold with totally geodesic boundary, its universal cover $\widetilde{N}$ may be identified with a convex subset of $\mathbb{H}^{3}$ bounded by a collection of geodesic hyperplanes. We will do so, and we will also continue to use the following terminology which originated in [26] and [27] and was used in [14].
Definition 2.1.1.
Let $N$ be a hyperbolic $3$-manifold with compact totally geodesic boundary, and let $\widetilde{N}\subset\mathbb{H}^{3}$ be its universal cover. A short cut in $\widetilde{N}$ is a geodesic arc joining the closest points of two distinct components of $\partial\widetilde{N}$. A return path, or orthogeodesic, in $N$ is the projection of a short cut under the universal covering map.
Each return path is a homotopically non-trivial geodesic arc properly immersed in $N$, perpendicular to $\partial N$ at each of its endpoints. Corollary 3.3 of [26] asserts that for a fixed $K\in\mathbb{R}$ and finite-volume hyperbolic manifold $N$ with totally geodesic boundary, there are only finitely many return paths in $N$ with length less than $K$. This makes possible the following:
Definition 2.1.2.
Let $N$ be a finite-volume hyperbolic $3$-manifold with compact totally geodesic boundary. Upon enumerating the collection of return paths as $\{\lambda_{1},\lambda_{2},\ldots\}$, where for each $i\in\mathbb{N}$ the length of $\lambda_{i+1}$ is at least the length of $\lambda_{i}$, let $\ell_{i}$ denote the length of $\lambda_{i}$. The orthospectrum of $N$ is the sequence $(\ell_{1},\ell_{2},\ldots)$. Its elements are ortholengths.
For $N$ as above, hyperbolic trigonometry relates the orthospectrum to arc lengths on $\partial N$ by means of a class of totally geodesic hexagons in $\widetilde{N}$ that have short cuts as some edges. Below we reproduce two lemmas from [14] that describe the hexagons in question.
Lemma 2.1.3 ([14], Lemma 2.3).
Suppose that $\Pi_{1}$, $\Pi_{2}$, and $\Pi_{3}$ are mutually disjoint geodesic planes in $\mathbb{H}^{3}$. For each two-element subset $\{i,j\}$ of $\{1,2,3\}$, let $\lambda_{ij}$ denote the common perpendicular to $\Pi_{i}$ and $\Pi_{j}$. Then $\lambda_{12}$, $\lambda_{13}$, and $\lambda_{23}$ lie in a common plane $\Pi$.
Lemma 2.1.4 ([14], Lemma 2.4).
Let $N$ be a finite-volume hyperbolic $3$-manifold with compact totally geodesic boundary, and suppose $\Pi_{1}$, $\Pi_{2}$, and $\Pi_{3}$ are distinct components of $\partial\widetilde{N}$. Let $\Pi$ be the plane, produced by Lemma 2.1.3, which contains the short cuts $\lambda_{12}$, $\lambda_{13}$, and $\lambda_{23}$. Let $C$ be the right-angled hexagon in $\Pi$ with edges the $\lambda_{ij}$ for $1\leq i<j\leq 3$, together with the set of geodesic arcs in the $\Pi_{i}$ joining their endpoints. Then $C\subset\widetilde{N}$, and $C\cap\partial\widetilde{N}=\bigcup_{i}(C\cap\Pi_{i})$.
In [14] the result above is stated only for compact hyperbolic $3$-manifolds with totally geodesic boundary, but its proof carries over to the current context without revision.
Definition 2.1.5.
For a finite-volume hyperbolic $3$-manifold NN with compact totally geodesic boundary and three components $\Pi_{1}$, $\Pi_{2}$, and $\Pi_{3}$ of $\widetilde{N}$, let $C$ be the right-angled hexagon supplied by Lemma 2.1.4. We call the edges of $C$ which are short cuts internal, and the edges in $\partial\widetilde{N}$ external. If the internal edges lift $\lambda_{i}$, $\lambda_{j}$, and $\lambda_{k}$, we call $C$ an $(i,j,k)$ hexagon.
We will say that the feet of a return path $\lambda$ are the points $\lambda\cap\partial N$, and similarly for the feet of a short cut. The orthospectrum of $N$ is related to the set of lengths of arcs in $\partial N$ joining feet of return paths.
Definition 2.1.6.
Let $N$ be a finite-volume hyperbolic $3$-manifold with compact totally geodesic boundary. For $i,j\in\mathbb{N}$ let $d_{ij}$ be the length of the shortest non-constant geodesic arc joining a foot of $\lambda_{i}$ to one of $\lambda_{j}$, or $\infty$ if no such arc exists. For any $k\in\mathbb{N}$ let $d_{ij}^{(k)}$ be the length of the $k$th-shortest such arc, or $\infty$ as appropriate.
Lemma 2.1.7.
For $i$, $j$, and $k$ in $\mathbb{N}$, let $X_{ij}^{k}$ be determined by
(2.1.7.1) $\cosh X_{ij}^{k}=\frac{\cosh\ell_{i}\cosh\ell_{j}+\cosh\ell_{k}}{\sinh\ell_{i}\sinh\ell_{j}}$
For a finite-volume hyperbolic $3$-manifold NN with compact totally geodesic boundary and any fixed $i,j\in\mathbb{N}$, if $k_{1}\leq k_{2}\leq k_{3}\ldots$ is the set of $k\in\mathbb{N}$ such that there exists an $(i,j,k)$ hexagon in $\widetilde{N}$ then $d_{ij}=X_{ij}^{k_{1}}$, and $d_{ij}^{(n)}=X_{ij}^{k_{n}}$ for $n>1$. In particular, $d_{ij}^{k}\geq X_{ij}^{k}$ for all $i,j,k\in\mathbb{N}$.
Proof.
Any given geodesic arc $\gamma$ on $\partial N$ that joins a foot of $\lambda_{i}$ to one of $\lambda_{j}$ lifts to a geodesic arc $\tilde{\gamma}$ on a component $\Pi$ of $\partial\widetilde{N}$ joining the foot of a lift $\tilde{\lambda}_{i}$ of $\lambda_{i}$ to that of a lift $\tilde{\lambda}_{j}$ of $\lambda_{j}$. The feet of $\tilde{\lambda}_{i}$ and $\tilde{\lambda}_{j}$ opposite their intersections with $\tilde{\gamma}$ lie in components of $\partial\widetilde{N}$ joined by a short cut $\tilde{\lambda}_{k}$ for some $k$. Lemma 2.1.4 then implies that $\tilde{\gamma}$ is the external edge of an $(i,j,k)$ hexagon opposite $\tilde{\lambda}_{k}$. The “right-angled hexagon rule” [31, Theorem 3.5.13] implies that the length of $\tilde{\gamma}$, hence of $\gamma$, is $X_{ij}^{k}$.
It is a quick consequence of the definition that for fixed $i$ and $j$, if $k<k^{\prime}$ then $X_{ij}^{k}<X_{ij}^{k^{\prime}}$. The lemma follows.∎
We close this subsection with some basic observations on the monotonicity of the $X_{ij}^{k}$ for $i,j,k\in\{1,2\}$.
Lemma 2.1.8.
The function $X_{11}^{1}$ defined Lemma 2.1.7 is decreasing in $\ell_{1}$. $X_{11}^{2}$ decreases in $\ell_{1}$ and increases in $\ell_{2}$, and $X_{12}^{1}$ decreases in both $\ell_{1}$ and $\ell_{2}$. Moreover, $X_{11}^{1}\geq X_{12}^{1}\geq X_{22}^{1}$.
Proof.
Each of these follows from (2.1.7.1) with a little manipulation. We have for instance:
(2.1.8.1) $\cosh X_{11}^{1}=\frac{\cosh^{2}\ell_{1}+\cosh\ell_{1}}{\sinh^{2}\ell_{1}}=\frac{\cosh\ell_{1}}{\cosh\ell_{1}-1}=1+\frac{1}{\cosh\ell_{1}-1},$
and the first assertion is clear. Along similar lines:
(2.1.8.2) $\cosh X_{11}^{2}=\frac{\cosh^{2}\ell_{1}+\cosh\ell_{2}}{\sinh^{2}\ell_{1}}=1+\frac{1+\cosh\ell_{2}}{\sinh^{2}\ell_{1}},$
and the second assertion holds. Finally:
(2.1.8.3) $\cosh X_{12}^{1}=\frac{\cosh\ell_{1}}{\sinh\ell_{1}}\sqrt{\frac{\cosh\ell_{2}+1}{\cosh\ell_{2}-1}}=\coth\ell_{1}\sqrt{1+\frac{2}{\cosh\ell_{2}-1}}$
That $X_{12}^{1}$ decreases in both $\ell_{1}$ and $\ell_{2}$ now follows from the fact that $\coth x$ decreases in $x$. Moreover, both $\cosh X_{11}^{1}$ and $\cosh X_{12}^{1}$ are of the form
$\frac{\cosh x\cosh\ell_{1}+\cosh\ell_{1}}{\sinh x\sinh\ell_{1}}=\coth x\coth\ell_{1}+\frac{\cosh\ell_{1}}{\sinh x\sinh\ell_{1}},$
where one substitutes $\ell_{1}$ for $x$ to produce $\cosh X_{11}^{1}$ and $\ell_{2}$ for $x$ to produce $\cosh X_{12}^{1}$. Since $x\mapsto\coth x$ is decreasing and $x\mapsto\sinh x$ is increasing, it follows that $\cosh X_{11}^{1}\geq\cosh X_{12}^{1}$. This implies the left-hand inequality above; the right-hand inequality follows similarly.∎
2.2. Kojima–Miyamoto’s lower bound
Here we will reproduce an argument originally from [27] which was slightly improved in [14], and improve it slightly further. It gives a lower bound on $\ell_{2}$ as a function of $\ell_{1}$ among finite-volume hyperbolic $3$-manifolds with compact totally geodesic boundary. This bound is rarely near sharp, but it is easily computable. The functions $R$, $R^{\prime}$, and $E$ of $\ell_{1}$ below are as in [27] and [14]; $R^{\prime\prime}$ and $M$ match [14].
The basic idea here is that for a hyperbolic $3$-manifold NN with compact totally geodesic boundary, the topology of $\partial N$ determines its area by the Gauss–Bonnet theorem, and this bounds the areas of disks in a packing of $\partial N$. The radii of such a packing are determined by the orthospectrum of $N$.
Lemma 2.2.1.
For $X_{ij}^{k}$ as defined in (2.1.7.1), let $R=X_{11}^{1}/2$, satisfying
(2.2.1.1) $\cosh R=\sqrt{\frac{2\cosh\ell_{1}-1}{2\cosh\ell_{1}-2}}=\sqrt{1+\frac{1}{2\cosh\ell_{1}-2}},$
and let $S=X_{12}^{1}-R$. For a finite-volume hyperbolic $3$-manifold NN with compact, connected totally geodesic boundary such that $S>0$ there are four disks embedded in $\partial N$ without overlapping: two of radius $R$, centered at feet of $\lambda_{1}$, and two of radius $S$, centered at feet of $\lambda_{2}$.
Proof.
By Definition 2.1.6, disks $U$ and $U^{\prime}$ of radius $r$, centered at the feet of $\lambda_{1}$, are embedded in $\partial N$ without overlapping each other if and only $r\leq d_{11}/2$. Similarly, disks $V$ and $V^{\prime}$ of radius $s$, centered at the feet of $\lambda_{2}$, are embedded in $\partial N$ without overlapping if and only if $s\leq d_{22}/2$. Finally, if $r+s\leq d_{12}$ then $U\cup U^{\prime}$ does not overlap $V\cup V^{\prime}$.
Lemma 2.1.7 and the definitions above imply that $R\leq d_{11}/2$ and $R+S\leq d_{12}$. We will show below that $S\leq X_{22}^{1}/2\leq d_{22}/2$, hence by the paragraph above that the lemma holds. Applying (2.1.7.1) and the “angle addition formula” for hyperbolic sine yields:
(2.2.1.2) $\sinh S=\frac{\sqrt{(2\cosh^{2}\ell_{1}+\cosh\ell_{2}-1)(2\cosh\ell_{1}-1)}-\cosh\ell_{1}\sqrt{\cosh\ell_{2}+1}}{(\cosh\ell_{1}-1)\sqrt{2(\cosh\ell_{1}+1)(\cosh\ell_{2}-1)}}$
$\sinh\left(X_{22}^{1}/2\right)=\frac{\sqrt{\cosh\ell_{1}+1}}{\sqrt{2}\sinh\ell_{2}}$
Subtracting $\sinh S$ from $\sinh\left(X_{22}^{1}/2\right)$ and using the common denominator $\sqrt{2}\sinh\ell_{2}(\cosh\ell_{1}-1)\sqrt{\cosh\ell_{1}+1}$, we find that the numerator of $\sinh\left(X_{22}^{1}/2\right)-\sinh S$ is as below:
$\cosh^{2}\ell_{1}-1+\cosh\ell_{1}(\cosh\ell_{2}+1)-2\sqrt{\left(\cosh^{2}\ell_{1}+\frac{\cosh\ell_{2}-1}{2}\right)\left(\cosh\ell_{1}-\frac{1}{2}\right)(\cosh\ell_{2}+1)}$
$=\ \ \left(\sqrt{\cosh^{2}\ell_{1}+\frac{\cosh\ell_{2}-1}{2}}-\sqrt{\left(\cosh\ell_{1}-\frac{1}{2}\right)(\cosh\ell_{2}+1)}\right)^{2}$
Therefore $\sinh\left(X_{22}^{1}/2\right)-\sinh S\geq 0$. Setting the above equal to zero and solving the resulting equation, we obtain $\cosh\ell_{1}=\cosh\ell_{2}$.∎
The main result of this subsection is the following improvement on Lemma 2.9 of [14]. In introducing it we recall that Kojima–Miyamoto proved that $\cosh\ell_{1}\geq\frac{3+\sqrt{3}}{4}$ for every compact hyperbolic $3$-manifold with connected, totally geodesic boundary of genus two [27, Corollary 3.5]. Their proof again carries through to the current setting.
Proposition 2.2.2.
For $R$ as in Lemma 2.2.1 let $R^{\prime}$ satisfy $\cosh R^{\prime}=3-\cosh R$ and define a function $E$ of $\ell_{1}$ by:
(2.2.2.1) $\cosh E=1+\frac{2}{\cosh^{2}(R+R^{\prime})\cdot\tanh^{2}\ell_{1}-1}$
$E$ is decreasing for $\frac{3+\sqrt{3}}{4}\leq\cosh\ell_{1}\leq 1.4$. For $R^{\prime\prime}$ determined by $\cosh R^{\prime\prime}=\frac{1}{2\sin(\pi/9)}=1.4619$..., define a quantity $M$ that depends on $\ell_{1}$ by
(2.2.2.2) $\cosh M=\sqrt{1+\frac{\cosh\ell_{1}+1}{\cosh(2R^{\prime\prime})-1}}$
For a finite-volume hyperbolic $3$-manifold NN with compact, connected totally geodesic boundary of genus $2$, $\ell_{2}\geq\max\{\ell_{1},E,M\}$.
Remark 2.2.3.
This combines and improves Lemmas 2.8 and 2.9 of [14], of which Lemma 2.8 re-recorded arguments in [27] (in and around Lemmas 4.2 and 4.3 there), and Lemma 2.9 gave a new, improved bound in a subinterval. Of the functions $E$, $F$, $L$, and $M$ there, the work in Lemma 2.2.1 above implies that $F\geq E$, and we argue directly below that $L\geq M$.
Proof.
If disks $U$ and $U^{\prime}$, of radius $R$, and disks $V$ and $V^{\prime}$ of radius $S$ are all embedded in $\partial N$ without overlapping then the sum of their areas is less than the area of $\partial N$, which is $4\pi$ by the Gauss–Bonnet theorem. We obtain the following inequality:
$4\pi(\cosh R-1)+4\pi(\cosh S-1)\leq 4\pi,\quad\Rightarrow\quad\cosh S\leq 3-\cosh R$
Thus for $R^{\prime}$ as defined above, $S\leq R^{\prime}$. Now taking $R$ and $S=X_{12}^{1}-R$ as defined in Lemma 2.2.1, and applying that result’s conclusion, we find that $X_{12}^{1}\leq R+R^{\prime}$. Since $X_{12}^{1}$ is decreasing in $\ell_{2}$ (recall Lemma 2.1.8), the upper bound on $X_{12}^{1}$ determines a lower bound on $\cosh\ell_{2}$. Setting $X_{12}^{1}$ equal to $R+R^{\prime}$ and solving for $\cosh\ell_{2}$ yields formula (2.2.2.1) for $\cosh E$. It was proved in [14, Lemma 3.4] that $E$, so defined, decreases for $\frac{3+\sqrt{3}}{4}\leq\cosh\ell_{1}\leq 1.4$.
It was proved in [14, Lemma 2.9] that $\ell_{2}$ is also bounded below by $\min\{L,M\}$, for $\ell_{1}$ satisfying:
$\cosh\ell_{1}\leq\frac{\cos(2\pi/9)}{2\cos(2\pi/9)-1}=1.43969$...
where $L$ and $M$ are respectively defined as functions of $\ell_{1}$ in formulas (2.9.2) and (2.9.3) there. (The formula for $M$ is reproduced above in (2.2.2.2).) We need only observe that $L$ is decreasing and $M$ is increasing as functions of $\ell_{1}$, and they agree at the right endpoint $\cos(2\pi/9)/(2\cos(2\pi/9)-1)$ of the relevant interval, to conclude that $\min\{L,M\}=M$ here.
Substituting $\cosh\ell_{1}=\cos(2\pi/9)/(2\cos(2\pi/9)-1)$ in (2.2.2.2), then simplifying, shows that $M=\ell_{1}$ here. For $\ell_{1}$ larger than this value, $M<\ell_{1}$, and it is true by definition that $\ell_{2}\geq\ell_{1}$, so on this interval we also have $\ell_{2}\geq\max\{\ell_{1},E,M\}$.∎
Remark 2.2.4.
It is worth noting just how far from sharp the lower bound on $\ell_{2}$ given by $E$ is at $\cosh\ell_{1}=\frac{3+\sqrt{3}}{4}$. For $N$ with this value of $\ell_{1}$, $\partial N$ decomposes into equilateral triangles of side length $\cosh^{-1}(3+2\sqrt{3})$, and an explicit calculation gives $\cosh\ell_{2}=\frac{13+9\sqrt{3}}{4}\simeq 7.147$. On the other hand, $\cosh E\simeq 2.893$ here.
2.3. A bound for $\ell_{1}$ in the absence of a $(1,1,1)$-hexagon
Proposition 3.9 of [14] bounds the first ortholength $\ell_{1}$ of a finite-volume hyperbolic $3$-manifold NN with compact totally geodesic boundary below by $\cosh\ell_{1}\geq 1.215$, assuming $\widetilde{N}$ has no $(1,1,1)$-hexagon. This is significantly better than the sharp universal lower bound of $\frac{3+\sqrt{3}}{4}\simeq 1.183$ for $\cosh\ell_{1}$ proved in [27]. (For comparison, all manifolds with volume less than $7.63$ in the census of Petronio et. al. have $\cosh\ell_{1}<1.213$.)
The main observation behind this result is simply that in the absence of a $(1,1,1)$-hexagon, the arc length $d_{11}$ on $\partial N$ is bounded below by $X_{11}^{2}$ instead of $X_{11}^{1}$; this follows from the first assertion of Lemma 2.1.7. Here we will recast this observation to give an upper bound on $\ell_{2}$ in terms of $\ell_{1}$, and also improve the absolute lower bound of $1.215$ for $\cosh\ell_{1}$, in this setting.
Proposition 2.3.1.
Let $N$ be a finite-volume hyperbolic $3$-manifold with compact, connected, totally geodesic boundary of genus two. If $\widetilde{N}$ has no $(1,1,1)$-hexagon then
$\cosh\ell_{2}\leq(2+2\sqrt{3})\sinh^{2}\ell_{1}-1,$
and $\cosh\ell_{1}>1.23$.
Proof.
It is a consequence of Boröczky’s theorem recorded in Corollary 3.5 of [27] and Lemma 2.7 of [14] that $\cosh d_{11}\leq 3+2\sqrt{3}$ with the hypotheses above. (As usual, this is stated in [14] for $N$ compact, but it extends without revision to the finite-volume case.) Using the fact that $d_{11}$ is at least $X_{11}^{2}$ in the absence of $(1,1,1)$-hexagons, substituting the right-hand side of formula (2.1.8.2) for $\cosh d_{11}$ in the inequality above, and solving for $\cosh\ell_{2}$ gives the upper bound on $\ell_{2}$ in terms of $\ell_{1}$.
For the absolute lower bound on $\cosh\ell_{1}$, we note that the upper bound we have just proved for $\ell_{2}$ is an increasing function of $\ell_{1}$, whereas the lower bound $E$ of Proposition 2.2.2 is decreasing. Direct computation shows that $E$ takes the value $1.200$... and the upper bound the value $1.194$... when $\cosh\ell_{1}=1.23$ (and their values coincide when $\cosh\ell_{1}\approx 1.2304$). Thus if there is no $(1,1,1)$-hexagon and $\cosh\ell_{1}\leq 1.23$, the lower bound on $\ell_{2}$ exceeds the upper bound, a contradiction.∎
3. More Better Muffins
A “muffin”, so named by Kojima and Miyamoto in [27] (see the discussion below Proposition 3.1 of that paper), is a member of a certain class of hyperbolic solids of rotation, shown in [27] to embed in a compact hyperbolic three-manifold with totally geodesic boundary and used in the main volume bound of [27]. We used the same class of muffins from [27] in [14], denoting them as $\mathrm{Muf}_{\ell_{1}}$” in Definition 3.1 there. As this notation suggests, members of this class are determined up to isometry by a single parameter, which in their application to volume bounds is the length $\ell_{1}$ of the shortest return path. Here we will continue to use the notation of [14] for this class of muffins (see Definitions and Remarks 3.1.6 below).
This section introduces a more general class of muffins, still hyperbolic solids of rotation, but now depending on two parameters $\ell$ and $R$ which are side lengths of certain reflectively symmetric hyperbolic pentagons. In Section 4 we will use muffins together with collars to give lower bounds on volume for a hyperbolic $3$-manifold NN with totally geodesic boundary satisfying certain bounds on the lengths of $\ell_{1}$ and $\ell_{2}$. Here we formally define these objects and establish technical results — for muffins in Section 3.1 and collars in Section 3.2 — that will allow us to show that they are disjointly embedded in $N$.
3.1. Muffins
Given $\ell,R>0$, let $Q$ be a hyperbolic Lambert quadrilateral—one with three right angles—such that its sides having right angles at both endpoints are of lengths $\ell/2$ and $R$. Doubling $Q$ across its edge $\omega$ opposite the one $\rho$ with length $R$ yields a reflectively symmetric hyperbolic pentagon $P$ with four right angles. The base $\lambda$ of PP—the side opposite the non-right vertex—has length $\ell$, and each of the two sides intersecting it have length $R$. A muffin is the solid in $\mathbb{H}^{3}$ obtained by rotating $P$ around $\lambda$. We illustrate this construction in Figure 3.1 and formalize its definition below.
[…]
Figure 3.1. Making a muffin.
Definition 3.1.1.
The muffin $M(\ell,R)$ with height $\ell$ and cap radius $R$ is the solid that results from rotating a reflectively symmetric pentagon $P$ with four right angles about its base $\lambda$ (the side opposite its not-necessarily-right angle, labeled in Figure 3.1), where $\ell$ and $R$ are respectively the lengths of $\lambda$ and of the sides of $P$ adjacent to $\lambda$—the side of $\rho$ of $Q$ in Figure 3.1, and its mirror image. We say that $M(\ell,R)$ is centered at $\lambda$. Its caps are the disks of radius $R$ obtained by rotating $\rho$ around its vertex $\rho\cap\lambda$ and likewise for its mirror image.
We further say that the muffin has waist radius $W$, the length of the side $\omega$ of $Q$ in Figure 3.1, and side altitude $A$, where $A$ is the length of the final side of $Q$. These lengths are related by the “quadrilateral rule” of hyperbolic trigonometry as follows:
(3.1.1.1) $\tanh A=\cosh R\tanh(\ell/2)$ $\tanh W=\cosh(\ell/2)\tanh R$
It follows that $M(\ell,R)$ is determined up to isometry by $\ell$ and $R$, or by $\ell$ and $W$.
Definitions and Remarks 3.1.2.
For any $n\geq 2$ and any totally geodesic subspace $\Pi\subset\mathbb{H}^{n}$, there is a retraction $\pi\colon\thinspace\mathbb{H}^{n}\to\Pi$ that sends each point of $\mathbb{H}^{n}$ to its unique nearest point in $\Pi$. (See [6], particularly Lemma 3.2 there.) We call this map the orthogonal projection to $\Pi$. It has the property that for each $x\in\Pi$, $\pi^{-1}(x)\subset\mathbb{H}^{n}$ is a totally geodesic subspace of complementary dimension to $\Pi$, intersecting $\Pi$ orthogonally in $\{x\}$.
We give an explicit formula for the orthogonal projection to a certain plane $\Pi\subset\mathbb{H}^{3}$ in the proof of Proposition 6.2.3.
Each assertion about muffins recorded in the Fact below follows from their construction as solids of rotation by an exercise in hyperbolic geometry.
Fact 3.1.3.
For any disjoint pair of totally geodesic planes $\Pi_{1},\Pi_{2}\subset\mathbb{H}^{3}$ that attain a minimum distance $\ell>0$, and any $R>0$, the copy $M$ of $M(\ell,R)$ centered on the shortest geodesic arc joining $\Pi_{1}$ to $\Pi_{2}$ is contained in the convex set bounded by $\Pi_{1}$, and $\Pi_{2}$ and intersects $\Pi_{1}\cup\Pi_{2}$ precisely in the union of its caps. For such a muffin $M$:
  • Every point of $M$ is at a distance at most $W$ from $\lambda$, where $\lambda$ and $W$ are respectively the center and waist radius of $M$. The interior of $M$ is contained in the open $W$-neighborhood of $\lambda$.
  • For $i=1$ or $2$, taking $C_{i}$ to be the cap of $M$ contained in $\Pi_{i}$ and $\pi_{i}\colon\thinspace\mathbb{H}^{3}\to\mathbb{H}^{3}$ to be the orthogonal projection to $\Pi_{i}$, we have $M\subset\pi_{i}^{-1}(C_{i})$.
The proof of the following fact about right-angled hexagons uses the case $n=2$ of orthogonal projection: for any hyperbolic geodesic $\lambda\subset\mathbb{H}^{2}$ we have an orthogonal projection $\pi:{\mathbb{H}}^{2}\to\lambda$, and $\pi^{-1}(x)\subset\mathbb{H}^{2}$ is also a geodesic for any $x\in\lambda$.
Proposition 3.1.4.
For any right-angled hexagon $C$ in $\mathbb{H}^{2}$, and any pair of opposite sides $\lambda$, $\lambda^{\prime}$ of $C$, there is a geodesic arc $\delta\subset C$ that meets each of $\lambda$ and $\lambda^{\prime}$ at right angles at a point in its interior.
Proof.
Let $\pi$ denote the orthogonal projection of $\mathbb{H}^{2}$ to the geodesic containing $\lambda$.
Because $C$ is the intersection of half-planes bounded by the geodesics containing its sides, and the edges of $C$ that share endpoints with $\lambda$ are orthogonal to $\lambda$, we have $\pi(C)=\lambda$. Hence $\mu\doteq\rho(\lambda^{\prime})$ is a subsegment of $\lambda$. For each point $P$ of $\mu$, the hyperbolic geodesic $\pi^{-1}(P)$ contains a segment $s_{P}\subset C$ which meets $\lambda$ perpendicularly at $P$ and has its other endpoint in $\lambda^{\prime}$. Let us now choose an endpoint $P_{0}$ of $\lambda^{\prime}$, let $c$ denote the length of $\mu$, and for each $x\in(0,c]$ let $P_{x}$ denote the point of $\mu$ whose distance from $P_{0}$ is $x$. For each $x\in[0,c]$ we set $s_{x}=s_{P_{x}}$. For every $x\in(0,c]$ there is a hyperbolic quadrilateral $Q_{x}\subset C$ whose sides are $s_{0}$, $s_{x}$, and subsegments $\mu_{x}$ and $\lambda^{\prime}_{x}$ of $\mu$ and $\lambda^{\prime}$ respectively. The two interior angles of $Q_{x}$ incident to $\mu_{x}$ are right angles. The interior angle between $\lambda^{\prime}_{x}$ and $s_{0}$ is independent of $x$, and will be denoted $\beta$, while the interior angle between $\lambda^{\prime}_{x}$ and $s_{x}$ will be denoted $\alpha(x)$. The angles $\beta$ and $\alpha(c)$ are each less than $\pi/2$, since they are sub-angles of vertex angles of $C$.
The side $\mu_{x}$ of $Q_{x}$ has length $x$; we will denote the length of its side $s_{0}$ by $a$. The hyperbolic law of cosines for quadrilaterals with two right angles, which is stated in [15, VI.3.3], gives:
$\cos\alpha(x)=-\cos\beta\cosh x+\sin\beta\sinh x\sinh a.$
Noting that $\cos\alpha(x)\to-\cos\beta$ as $x\to 0$, so that $\alpha(x)\to\pi-\beta>\pi/2$, and recalling that $\alpha(c)<\pi/2$, we conclude from the intermediate value theorem that $\alpha(x_{0})=\pi/2$ for some $x_{0}\in(0,c^{\prime})$. Hence the segment $\delta\doteq s_{x_{0}}\subset C$ is perpendicular to both $\lambda$ and $\lambda^{\prime}$. ∎
For a given hyperbolic manifold $N$ with totally geodesic boundary, we will embed a muffin in NN—for carefully chosen $\ell$ and RR—by checking that a copy of $M(\ell,R)$ carefully placed in the universal cover $\widetilde{N}$ does not intersect its translates under the action of $\pi_{1}(N)$ by covering transformations. Here as in the prior works [27] and [14], “carefully placed” will always mean centered at a lift $\tilde{\lambda}$ of a return path $\lambda$, with height $\ell$ equal to the length of $\lambda$ so that each of its caps lies in a component of $\partial\widetilde{N}$. We use the following result to ensure that $M(\ell,R)$, with choices carefully made, lies in $\widetilde{N}$.
Lemma 3.1.5.
Let $N$ be a hyperbolic manifold with geodesic boundary, and let $\widetilde{N}$ denote its universal cover. If $\tilde{\lambda}_{k}$ is a lift of the $k$th-shortest return path to $\widetilde{N}$, the distance from $\tilde{\lambda}_{k}$ to any component of $\partial\widetilde{N}$ that does not contain either of its endpoints is at least $U_{k}$ defined by $\sinh U_{k}=\cosh\ell_{1}/\sinh(\ell_{k}/2)$. Equality is attained if $\tilde{\lambda}_{k}$ is an edge of a $(1,1,k)$-hexagon in $\widetilde{N}$.
Proof.
For $\tilde{\lambda}_{k}$ as above, let $\Pi_{1}$ and $\Pi_{2}$ be the components of $\partial\widetilde{N}$ containing its endpoints. For a third component $\Pi_{3}$ of $\widetilde{N}$, let $C$ be the planar right-angled hexagon supplied by Lemma 2.1.4, containing $\tilde{\lambda}_{k}$ and intersecting each $\Pi_{i}$ perpendicularly in a side, for $i=1$, 22, 33. Then $C\cap\Pi_{3}$ is the side of $C$ opposite $\tilde{\lambda}_{k}$.
We now apply Proposition 3.1.4, with $\tilde{\lambda}_{k}$ and $C\cap\Pi_{3}$ playing the roles of $\lambda$ and $\lambda^{\prime}$ in that proposition. Let $\delta\subset C$ be the arc supplied by Proposition 3.1.4. The hexagon $C$ is divided into two right-angled pentagons by $\delta$. Let us take the length of $\delta$ to be $y$ and $\ell_{i}$, $\ell_{j}$, $\ell_{k}$ the lengths of $\tilde{\lambda}_{i}$, $\tilde{\lambda}_{j}$, and $\tilde{\lambda}_{k}$, respectively, and $x_{i}$ and $x_{j}$ those of the sub-arcs of $\tilde{\lambda}_{k}$ in the pentagons containing $\tilde{\lambda}_{i}$ and $\tilde{\lambda}_{j}$. So $x_{i}+x_{j}=\ell_{k}$, and the “law of sines” for mostly-right-angled pentagons recorded in [15, VI.3.2] gives:
$\cosh\ell_{i}=\sinh x_{i}\sinh y\quad\mbox{and}\quad\cosh\ell_{j}=\sinh x_{j}\sinh y.$
Supposing without loss of generality that $\ell_{j}\geq\ell_{i}$, the equations above combine to imply that $x_{j}\geq x_{i}$ and hence that $x_{i}\leq\ell_{k}/2$. Using the first equation we now obtain:
$\sinh y=\frac{\cosh\ell_{i}}{\sinh x_{i}}\geq\frac{\cosh\ell_{1}}{\sinh(\ell_{k}/2)}$
The inequality above comes from the fact above that $x_{i}\leq\ell_{k}/2$, using the default bound $\ell_{i}\geq\ell_{1}$ in the numerator. Noting that if $C$ is a $(1,1,k)$ hexagon then $\delta$ is its axis of reflective symmetry and hence $x_{i}=x_{j}=\ell_{k}/2$, we obtain the Lemma’s final assertion about equality. ∎
We will use Lemma 3.1.5 to give a sufficient condition in Lemma 3.1.8 below, for a copy of $M(\ell,r)$ to lie entirely in $\widetilde{N}$. First we formally establish the link between the muffin $\mathrm{Muf}_{\ell_{1}}$ originally defined in [27] and [14, Dfn. 3.1], and $M(\ell,R)$ defined in this paper.
Definitions and Remarks 3.1.6.
In the notation of Definition 3.1.1, the muffin $\mathrm{Muf}_{\ell_{1}}$ originally defined in [27] and [14, Dfn. 3.1] is $M(\ell_{1},R(\ell_{1}))$, where $R=R(\ell_{1})$ is given by the formula (2.2.1.1). From (3.1.1.1) we therefore obtain the following formulas determining its side altitude $A=A(\ell_{1})$ and waist radius $W=W(\ell_{1})$ as functions of $\ell_{1}$ only:
(3.1.6.1) $\begin{split}\tanh A&=\cosh R\tanh(\ell_{1}/2)=\sqrt{\frac{2\cosh\ell_{1}-1}{2\cosh\ell_{1}+2}}\\ \tanh W&=\cosh(\ell_{1}/2)\tanh R=\sqrt{\frac{\cosh\ell_{1}+1}{4\cosh\ell_{1}-2}}\end{split}$
The geometric motivation for the choice of cap radius for $\mathrm{Muf}_{\ell_{1}}$ is captured in Figure 3.1 of [27]. The left side of that Figure pictures a $(1,1,1)$-hexagon in our notation, with $\ell$” there equal to $\ell_{1}$ and “RR” and “AA” given by (2.2.1.1) and (3.1.6.1). The Fact below records a feature of that picture.
Fact 3.1.7.
For $U_{1}$ as in Lemma 3.1.5 and $A=A(\ell_{1})$ and $W=W(\ell_{1})$ as in 3.1.6.1, we have $U_{1}=A+W$.
Proof.
Manipulating the formula that defines $U_{1}$ in Lemma 3.1.5 gives
$\tanh U_{1}=\frac{\cosh\ell_{1}}{\sqrt{\cosh^{2}\ell_{1}+\sinh^{2}(\ell_{1}/2)}}=\frac{\sqrt{2}\cosh\ell_{1}}{\sqrt{(2\cosh\ell_{1}-1)(\cosh\ell_{1}+1)}}$
The result now follows from the identity $\tanh(x+y)=\frac{\tanh x+\tanh y}{1+\tanh x\tanh y}$, using the formulas for the side altitude $A(\ell_{1})$ and waist radius $W(\ell_{1})$ of $\mathrm{Muf}_{\ell_{1}}$ from (3.1.6.1). ∎
We use Fact 3.1.7 in the proof of Lemma 3.1.8 immediately below, then in a stronger way in the proof of Proposition 3.2.2.
Lemma 3.1.8.
For a hyperbolic $3$-manifold NN and a lift $\tilde{\lambda}_{k}$ of the $k$th-shortest return path of $N$ to the universal cover $\widetilde{N}$, a copy of $M(\ell_{k},R)$ centered at $\tilde{\lambda}_{k}$ is contained in $\widetilde{N}$ if its waist radius is less than the quantity $U_{k}$ defined in Lemma 3.1.5. If so, it intersects $\partial\widetilde{N}$ precisely in its caps.
In particular, a copy of $\mathrm{Muf}_{\ell_{1}}$ centered at a lift $\tilde{\lambda}_{1}$ of $\lambda_{1}$ is contained in $\widetilde{N}$.
Proof of Lemma 3.1.8.
It is useful here to recall that $\widetilde{N}$ is an intersection of hyperbolic half-spaces in $\mathbb{H}^{3}$, each bounded by a totally geodesic plane that is a component of the preimage of $\partial N$. By construction of $M(\ell_{k},R)$, its caps are totally geodesic disks that each intersects its central geodesic arc at right angles at an endpoint. Since $\tilde{\lambda}_{k}$ intersects a component of $\partial\widetilde{N}$ at right angles at each of its endpoints, upon embedding $M(\ell_{k},R)$ in $\mathbb{H}^{3}$ so that its center coincides with $\tilde{\lambda}_{k}$, each cap is contained in one of these components, and $M(\ell_{k},R)$ itself is contained in the intersection of the two half-spaces that they bound containing $\widetilde{N}$.
If $M(\ell_{k},R)$ were not entirely contained in $\widetilde{N}$, it would thus intersect a third component of $\partial\widetilde{N}$, i.e. one not containing either cap. By Lemma 3.1.5, any such component has distance at least $U_{k}$ from $\tilde{\lambda}_{k}$. However, as observed in Fact 3.1.3, $M(\ell_{k},R)$ is contained in the $W$-neighborhood of its central geodesic arc, where $W$ is its waist radius. This implies that if $W<U_{k}$ then $M(\ell_{k},R)$, so embedded in $\mathbb{H}^{3}$, is entirely contained in $\widetilde{N}$ and intersects $\partial\widetilde{N}$ only in its caps.
By Fact 3.1.7 we have $U_{1}=A+W$. Therefore $W<U_{1}$, so by the above, a copy of $\mathrm{Muf}_{\ell_{1}}$ centered at a lift of $\lambda_{1}$ is entirely contained in $\widetilde{N}$. ∎
Kojima–Miyamoto proved in [27, Lemma 3.2], by an ad hoc argument, that a copy of $\mathrm{Muf}_{\ell_{1}}$ centered at a lift of $\lambda_{1}$ embeds in $N$. Here we develop a more systematic approach with the goal of identifying when two muffins, centered at lifts of $\lambda_{1}$ and $\lambda_{2}$, embed disjointly in $N$.
Here is the key observation. Let $N$ be a compact hyperbolic $3$-manifold with totally geodesic boundary, and let $M$ and $M^{\prime}$ be muffins centered at lifts $\tilde{\lambda}$ and $\tilde{\lambda}^{\prime}$ to $\widetilde{N}$ of return paths of $N$. Let $\Pi_{1}$ and $\Pi_{2}$ be the components of $\partial\widetilde{N}$ containing the caps of $M$, and let $\Pi_{1}^{\prime}$ and $\Pi_{2}^{\prime}$ play the same role for $M^{\prime}$. If $\Pi_{i}\neq\Pi_{j}^{\prime}$ for $i,j\in\{1,2\}$, then the four planes $\Pi_{1},\Pi_{2},\Pi_{1}^{\prime},\Pi_{2}^{\prime}$ determine a truncated tetrahedron $\Delta$, for which the distance from $\tilde{\lambda}$ to $\tilde{\lambda}^{\prime}$ is a transversal length. These notions are defined carefully in the preprint [13], which proves lower bounds on transversal length that we will use here to ensure that muffins do not overlap. Below we review their definitions in a form adapted to the current context.
Let $\mathcal{B}=\{\Pi_{1},\Pi_{2},\Pi_{1}^{\prime},\Pi_{2}^{\prime}\}$, and for $i=1,2$, let $H_{i}$ (respectively, $H_{i}^{\prime}$) be the half-space bounded by $\Pi_{i}$ (resp. $\Pi_{i}^{\prime}$) that contains $\widetilde{N}$ and hence the other $\Pi_{j}$ and $\Pi_{j}^{\prime}$. Recall from Lemma 2.1.3 that each three-element subset of $\mathcal{B}$ determines a geodesic plane that is perpendicular to all three of its members. It follows from Lemma 3.1 of [13] that either all four planes above have a common perpendicular plane $\Pi$, or that the common perpendicular to each three-plane subcollection bounds a single half-space that contains all three lifted return paths joining its members to the fourth plane. The truncated tetrahedron $\Delta$ determined by $\mathcal{B}$ is described in [13, Definition 3.2]. In the latter case, it is the intersection of the $H_{i}$ and $H_{i}^{\prime}$ with the four half-spaces determined in this way by the three-element subcollections of $\mathcal{B}$. In the former, we take $\Delta$ to be the intersection of the common perpendicular plane $\Pi$ with the $H_{i}$ and $H_{i}^{\prime}$, and say it is degenerate.
The internal edges of the truncated tetrahedron $\Delta$ defined as above are the lifted return paths joining each pair of distinct elements of $\mathcal{B}$. We say that two internal edges are opposite if no single member of $\mathcal{B}$ contains an endpoint of each. (Thus each internal edge is opposite a unique other internal edge.) The minimum distance between a pair of opposite internal edges is a transversal length of $\Delta$. (In general, its value depends on the choice of edges).
In the present setting, the lifts $\tilde{\lambda}$ and $\tilde{\lambda}^{\prime}$ are opposite edges of $\Delta$. Using results of [13], one can then bound the distance between $\tilde{\lambda}$ and $\tilde{\lambda}^{\prime}$ in terms of their lengths and the lengths of the other edges of $\Delta$. We illustrate the use of this philosophy below with a re-proof of [27, Lemma 3.2].
Lemma 3.1.9.
Let $N$ be an orientable hyperbolic $3$-manifold with compact totally geodesic boundary of genus $2$. For $\mathrm{Muf}_{\ell_{1}}$ as in Definitions and Remarks 3.1.6, a copy of $\mathrm{Muf}_{\ell_{1}}$ embedded in $\widetilde{N}$ centered at a lift of $\lambda_{1}$ is embedded in $N$ by the universal covering.
Proof.
By Lemma 3.1.8, a copy $M$ of $\mathrm{Muf}_{\ell_{1}}$ centered at a lift of $\lambda_{1}$ is contained in $\widetilde{N}$. The universal cover thus embeds it in $N$ if it is disjoint from all of its translates by the action of $\pi_{1}(N)$ on $\widetilde{N}$. These translates are copies of $\mathrm{Muf}_{\ell_{1}}$ centered at other lifts of $\lambda_{1}$. If a component $\Pi$ of $\partial\widetilde{N}$ contains a cap of $M$ and that of a translate $M^{\prime}$, then $M\cap M^{\prime}=\emptyset$. This is because each of $M$ and $M^{\prime}$ is contained in the preimage under orthogonal projection of its intersection with $\Pi$, and disks of radius $R$ about feet of lifts of $\lambda_{1}$ are disjoint or equal (as follows from Lemma 2.2.1).
Now suppose that $M$ and a translate $M^{\prime}$ do not have caps on a common component of $\partial\widetilde{N}$. Then the four planes containing their caps determine a truncated tetrahedron $\Delta$ as in [13, Definition 3.2]. Let $\lambda$ and $\lambda^{\prime}$ be the centers of $M$ and $M^{\prime}$, respectively, each a lift of $\lambda_{1}$, hence having length $\ell_{1}$. The distance from $\lambda$ to $\lambda^{\prime}$ is a transversal length of $\Delta$, again as in [13, Definition 3.2]. It is therefore given by $T(x,x;a,b,c,d)$ as in [13, Lemma 3.4], where $x=\cosh\ell_{1}$, and $a$, $b$, $c$, and $d$ are the hyperbolic cosines of lengths of the distinct edges of $\Delta$ not equal to $\lambda$ or $\lambda^{\prime}$. Since these other edges of $\Delta$ are also lifts of return paths, each also has length at least $\ell_{1}$. Therefore by [13, Proposition 3.5] we have:
$\cosh T(x,x;a,b,c,d)\geq\cosh D\doteq\frac{2x}{x-1}$
On the other hand, the waist radius $W$ of $M$ and $M^{\prime}$ is determined by (3.1.6.1). Since each of $M$ and $M^{\prime}$ is contained in the $W$-neighborhood of its center, $M$ does not intersect $M^{\prime}$ if $W<D/2$. Using the formula above for $D$, we obtain:
$\tanh(D/2)=\sqrt{\frac{x+1}{3x-1}}=\sqrt{\frac{\cosh\ell_{1}+1}{3\cosh\ell_{1}-1}}$
Since 3x1<4x23x-1<4x-2 for $x>1$ and $y\mapsto\tanh y$ is increasing, it follows that $D/2>W$, hence that $M$ does not intersect $M^{\prime}$. ∎
We now turn our attention to the problem of embedding two muffins, disjointly, centered at the feet of the shortest and second-shortest return paths. We will use the following helpful sign computation.
Lemma 3.1.10.
For $R$ as in (2.2.1.1) and $X_{ij}^{k}$ as in (2.1.7.1), regard $S=X_{12}^{1}-R$ as a function of variables $x=\cosh\ell_{1}$ and $y=\cosh\ell_{2}$. On the region defined by $x>1$ and $y\leq 3$, the function $S$ is strictly increasing in the variable $x$, strictly decreasing in the variable $y$, and positive-valued.
Proof.
We note that since $R$ depends only on $x$, it follows from Lemma 2.1.8 that $S$ is strictly decreasing in $y$.
We now claim that $\partial S/\partial x>0$. From (2.2.1.1) we obtain that $\sinh R=(2x-2)^{-1/2}$, hence:
$\frac{\partial R}{\partial x}=\frac{-(2x-2)^{-3/2}}{\cosh R}=\frac{-1}{(2x-2)\sqrt{2x-1}}$
Using the description of $X_{12}^{1}$ from (2.1.8.3), we obtain:
$\frac{\partial X_{12}^{1}}{\partial x}$ $=\frac{1}{\sinh X_{12}^{1}}\cdot\frac{-1}{(x^{2}-1)^{3/2}}\sqrt{\frac{y+1}{y-1}}=\frac{-1}{x^{2}-1}\sqrt{\frac{y+1}{y+2x^{2}-1}}$
Therefore
(3.1.10.1) $(x-1)\frac{\partial S}{\partial x}=\frac{1}{2\sqrt{2x-1}}-\frac{1}{x+1}\sqrt{\frac{y+1}{y+2x^{2}-1}}$
The right-hand side above vanishes at $x=1$, but we are interested only in $x>1$. Setting the right-hand side above equal to 0 and solving for $y$ yields
$y=\frac{2x^{3}+6x^{2}+7x-3}{5-x}$
This is an increasing function of $x$ with values greater than $3$ for $x>1$. Inserting the test values $x=2$ and $y=2$ into the right-hand side of (3.1.10.1) gives a positive value. Thus for $x$ and $y$ in the range of interest, $\partial S/\partial x(x,y)>0$, proving the claim. It therefore follows that $S$ is strictly increasing in $x$ for $x>1$ and $y\leq 3$.
We note that by (2.2.1.1) and (2.1.8.3), both $R$ and $X_{12}^{1}$ increase without bound for any fixed $y$ as $x$ approaches $1$ from the right. However, the formula for $\sinh S$ obtained in (2.2.1.2) shows that $S(x,3)$ has the limit $\sqrt{2}/4$ as $x$ approaches $1$ from above. In view of the monotonicity established above, it follows that $S$ is bounded below by $\sqrt{2}/4$, and is in particular strictly positive, for $x>1$ and $y\leq 3$. ∎
The first main result of this section is the muffin embedding criterion below. Based on numerical exploration, we expect that in general, muffins which are sized so that their caps embed in $\partial N$ without overlapping will themselves embed in $N$ without overlapping. But having not been able to prove this general expectation, we settle here for a criterion that can be numerically checked on a rectangle of possible values of $\ell_{1}$ and $\ell_{2}$.
Proposition 3.1.11.
Regard $R$ as in (2.2.1.1), $S=X_{12}^{1}-R$ as in Lemma 2.2.1, for $X_{ij}^{k}$ as in (2.1.7.1), and $W$ as in (3.1.6.1), as functions of variables $x=\cosh\ell_{1}$ and $y=\cosh\ell_{2}$. Suppose $N$ is an orientable hyperbolic $3$-manifold with $\partial N$ compact, connected, totally geodesic and of genus $2$, such that the lengths $\ell_{1}$ and $\ell_{2}$ of its shortest and second-shortest return paths satisfy
$a\leq\cosh\ell_{1}\leq b\quad\mbox{and}\quad c\leq\cosh\ell_{2}\leq d,$
for some given $a,b,c,d$ with $1<a<b\leq 2$ and $1<c<d\leq 3$. If $W_{S}^{0}\leq U_{2}^{0}$, $W_{R}^{0}+W_{S}^{0}\leq T_{12}^{0}$, and $2W_{S}^{0}\leq T_{22}^{0}$, where $\sinh U^{0}_{2}=\frac{a\sqrt{2}}{\sqrt{d-1}}$,
$W_{R}^{0}=\tanh^{-1}\left(\sqrt{\frac{a+1}{4a-2}}\right),$ $W_{S}^{0}=\tanh^{-1}\left(\sqrt{\frac{d+1}{2}}\tanh S(b,c)\right),$
(3.1.11.1) $T_{12}^{0}=\cosh^{-1}\left(\frac{2b}{\sqrt{(b-1)(d-1)}}\right),$ $\mbox{and}\quad T_{22}^{0}=\cosh^{-1}\left(\frac{2a}{d-1}\right),$
then the interior of a copy of $M(\ell_{2},S)$, centered at a lift of $\lambda_{2}$, is contained in $\widetilde{N}$ and embeds in $N$ without overlapping that of $\mathrm{Muf}_{\ell_{1}}$ from Lemma 3.1.9. (Note that by Lemma 3.1.10, the quantity $S(x,y)$ is positive-valued for $a\leq x\leq b$ and $c\leq y\leq d$, so that $W^{0}_{S}$ is well-defined and positive, and $M(\ell_{2},S)$ is defined.)
Proof.
We claim first that $W_{S}^{0}$ is an upper bound for the waist radius $W_{2}$ of $M(\ell_{2},S)$, taken as a function of $x$ and $y$, on the rectangle $a,b]\times[c,d]$. By the formula (3.1.1.1), $W_{2}$ satisfies $\tanh W_{2}=\cosh(\ell_{2}/2)\tanh S$. Lemma 3.1.10 implies that as a function of $x$ and $y$, $S$ increases with $x$ and decreases with $y$ at all points of this rectangle. The values of $S$ are thus bounded above by $S(b,c)$ on $a,b]\times[c,d]$, and the claim thus follows from the “half-angle identity” for hyperbolic cosine.
For $a\leq\cosh\ell_{1}\leq b$ and $c\leq\cosh\ell_{2}\leq d$, the constant $U_{2}^{0}$ defined in the statement of the proposition is a lower bound for the quantity $U_{2}$ defined in Lemma 3.1.5. Combining this observation with the hypothesis $W_{S}^{0}\leq U_{2}^{0}$ and the claim just proved, we deduce that on $W_{2}\leq U_{2}$. Lemma 3.1.8 then implies that a copy of $M(\ell_{2},S)$ centered at a lift of $\lambda_{2}$ is contained in $\widetilde{N}$.
To show that the interior of $M(\ell_{2},S)$ embeds in $N$ is equivalent to showing that it does not overlap any of its translates under the action of $\pi_{1}(N)$ on the universal cover $\widetilde{N}$ by covering transformations. Likewise, showing that the image of $M(\ell_{2},S)$ is disjoint in $N$ from that of $\mathrm{Muf}_{\ell_{1}}$ is equivalent to showing that $M(\ell_{2},S)$ does not overlap any translate of the latter muffin. We take these tasks on below.
Suppose first that a single component $\Pi$ of $\partial\widetilde{N}$ contains caps of both $M(\ell_{2},S)$ and either a translate of $\mathrm{Muf}_{\ell_{1}}$ or of itself. In the first case, these caps are discs of radius $S$ and $R$ in $\Pi$, respectively, centered at the feet of lifts of $\lambda_{2}$ and $\lambda_{1}$; in the second, both are of radius $S$ and centered at feet of lifts of $\lambda_{2}$. By Lemma 2.2.1, disks of radii $S$ and $R$ are disjointly embedded in $\partial N$ around the feet of $\lambda_{2}$ and $\lambda_{1}$; hence in either case the disks in question here do not overlap. Therefore, since each muffin is contained in the preimage of its cap under the orthogonal projection to $\Pi$, $M(\ell_{2},R)$ also does not overlap this translate of $\mathrm{Muf}_{\ell_{1}}$, or in the second case, of itself.
Now consider a translate of $\mathrm{Muf}_{\ell_{1}}$ or of $M(\ell_{2},S)$ with the property that neither of its caps is contained in a component of $\partial\widetilde{N}$ that also contains a cap of $M(\ell_{2},S)$. Let $\tilde{\lambda}_{2}$ be the lift of $\lambda_{2}$ at which $M(\ell_{2},S)$ is centered. In the first case let $\tilde{\lambda}_{1}$ be the center of the copy of $\mathrm{Muf}_{\ell_{1}}$ in question, and in the second let $\tilde{\lambda}_{2}^{\prime}$ be the center of the translate of $M(\ell_{2},S)$. There is a truncated tetrahedron $\Delta$ with $\tilde{\lambda}_{2}$ as one edge and either $\tilde{\lambda}_{1}$ or $\tilde{\lambda}_{2}^{\prime}$ its opposite, depending on the case, whose other edges are also lifts of return paths of $N$. Applying [13, Lemma 3.4] and [13, Proposition 3.5] as in the proof of Lemma 3.1.9 shows that the transversal length of $\Delta$ is bounded below by a function $T_{12}(x,y)$ or $T_{22}(x,y)$ in the respective cases, given by
$T_{12}(x,y)=\cosh^{-1}\left(\frac{2x}{\sqrt{(x-1)(y-1)}}\right)\quad\mbox{and}\quad T_{22}(x,y)=\cosh^{-1}\left(\frac{2x}{y-1}\right),$
for $x=\cosh\ell_{1}$ and $y=\cosh\ell_{2}$. Both $T_{12}$ and $T_{22}$ plainly decrease with $y$, for fixed $x$, and a computation shows that $\partial T_{12}/\partial x(x,y)<0$ for $x<2$ whereas $T_{22}$ increases with $x$. Therefore their values on the rectangle $a\leq x\leq b$, $c\leq y\leq d$ are respectively bounded below by $T^{0}_{12}=T_{12}(b,d)$ and $T_{22}^{0}=T_{22}(a,d)$, as given above in (3.1.11).
We have seen that the waist radius $W_{2}$ of $M(\ell_{2},S)$ is bounded above by $W^{0}_{S}$. We will similarly show below that the waist radius of $\mathrm{Muf}_{\ell_{1}}$ is bounded above by $W^{0}_{R}$. As recorded in Fact 3.1.3, a muffin with waist radius $W$ is contained in the $W$-neighborhood of its center. Thus the hypotheses $W^{0}_{R}+W^{0}_{S}\leq T_{12}^{0}$ and $2W^{0}_{S}\leq T_{22}^{0}$ will imply, respectively, that the interior of $M(\ell_{2},S)$ is disjoint from the translate of $\mathrm{Muf}_{\ell_{1}}$ and from $M(\ell_{2},S)$.
The waist radius of $\mathrm{Muf}_{\ell_{1}}$ is $W(x)$ as given by the formula (3.1.6.1). Manipulating that formula shows that $W$ is a decreasing function of $x$ for $x>1/2$, and hence is bounded above on the interval $a\leq x\leq b$ by $W(a)=W^{0}_{R}$, as asserted above. ∎
3.2. Collars
We now turn our attention to embedding collars, and controlling their interaction with muffins. The definition of collar given below is implicitly used in [27] and [14]. To motivate it, we recall from Section 2.1 that for a hyperbolic $3$-manifold NN with totally geodesic boundary, we take the universal cover $\widetilde{N}$ of $N$ to be a convex subset of $\mathbb{H}^{3}$ bounded by a collection of geodesic hyperplanes. For a component $\Pi$ of $\partial\widetilde{N}$, let $\pi$ be the orthogonal projection to $\Pi$ from Definitions and Remarks 3.1.2. Then for $x\in\Pi$, $\pi^{-1}(x)$ is a geodesic intersecting $\Pi$ orthogonally at $x$, and $\pi^{-1}(x)\cap\widetilde{N}$ is either a ray with endpoint $x$ or a segment with one endpoint at $x$ and the other on a different component of $\partial\widetilde{N}$.
Definition 3.2.1.
Let $N$ be a hyperbolic $3$-manifold with totally geodesic boundary. If $x$ is a point of $\partial N$ such that there exists a (necessarily unique) geodesic path in $N$ (parametrized by arclength) which begins at $x$, is perpendicular to $\partial N$ at its initial point, and has terminal point in $\partial N$, we shall denote the length of this path by $h_{x}$, the interval $0,h_{x}]$ by $J_{x}$, and the path itself by $\alpha_{x}:J_{x}\to N$. If $x$ is a point of $\partial N$ for which no path of this type exists, then there is a unique geodesic ray beginning at $x$; in this case, we shall set $h_{x}=+\infty$ and $J_{x}=[0,\infty)$, and denote (the arclength parametrization of) the ray by $\alpha_{x}:[0,\infty)\to N$. For a subset $S$ of $\partial N$, we define the open collar of $S$ in $N$ with height $h>0$ to be the set $\bigcup_{x\in S}\alpha_{x}([0,\min(h,h_{x})))$. We shall say that the collar is embedded if $h\leq\inf\{h_{x}\,|\,x\in S\}$, and the map from $S\times[0,h]$ to $N$ defined by $(x,t)\mapsto\alpha_{x}(t)$ is one-to-one.
We first state a result that was implicitly used, although not explicitly stated, in the main volume bound of [27] and [14].
Proposition 3.2.2.
Suppose $N$ is an orientable hyperbolic $3$-manifold with $\partial N$ compact, connected, totally geodesic and of genus $2$. Let $M$ be the projection to $N$ of a copy of the muffin $\mathrm{Muf}_{\ell_{1}}$ from Definitions and Remarks 3.1.6, centered at a lift of $\lambda_{1}$ to the universal cover $\widetilde{N}$, and let $C,C^{\prime}\subset\partial N$ be the two caps of $M$. For $A$ defined as in (3.1.6.1), an open collar of $\partial N-(C\cup C^{\prime})$ with any given height $h\leq\min\{A,\ell_{2}/2\}$ is embedded in $N$, disjointly from the interior of $M$.
Proof.
For a totally geodesic plane $\Pi\subset\mathbb{H}^{3}$, the “Hadamard–Hermann theorem” [22] implies that the exponential map restricts to a diffeomorphism from the normal bundle $\nu(\Pi)$ of $\Pi$ onto $\mathbb{H}^{3}$. Since an orientation of the bundle $\nu(\Pi)$ gives an identification of its total space with $\Pi\times\mathbb{R}$, it follows that there is a diffeomorphism from $\Pi\times(-h,h)$ to a metric open neighborhood of $\Pi$ in $\mathbb{H}^{3}$ taking $(x,t)$ to $\alpha_{x}(t)$ for any $x\in\Pi$ and any $t\in(-h,h)$, where $\alpha_{x}\colon\thinspace\mathbb{R}\to\mathbb{H}^{3}$ is an arclength-parametrized geodesic intersecting $\Pi$ orthogonally at $\alpha_{x}(0)=x$ for a choice of $\alpha_{x}^{\prime}(0)$ depending continuously on $x$.
Now if $N$ is a hyperbolic manifold with totally geodesic boundary as in the statement, we let $\widetilde{N}\subset{\mathbb{H}}^{3}$ denote the universal covering of $N$, we take $\Pi$ to be a component of $\partial\widetilde{N}$, we take $h\leq\min\{A,\ell_{2}/2\}$, and we choose the orientation of the bundle $\nu(\Pi)$ in such a way that $\alpha_{x}^{\prime}(0)$ points into $\widetilde{N}$ for each $x\in\Pi$. We claim that for each $x\in\Pi$ outside the preimage of $C\cup C^{\prime}$, and each $t\in[0,h)$, $\Pi$ is the unique closest component of $\partial\widetilde{N}$ to $\alpha_{x}(t)\in\widetilde{N}$. In particular, this implies that the quantity denoted $h_{x}$ in Definition 3.2.1 is at least $h$.
Suppose the claim does not hold, fix some $x_{0}\in\Pi$ outside the preimage of $C\cup C^{\prime}$ for which it fails, and let $t_{0}$ be the infimum of the set of $t\in[0,h)$ such that a component $\Pi^{\prime}\neq\Pi$ of $\partial\widetilde{N}$ is at least as close to $\alpha_{x_{0}}(t)$ as $\Pi$. Since $t_{0}$ is infimal with this property, $\alpha_{x_{0}}(t_{0})$ has distance $t_{0}$ from $\Pi^{\prime}$ as well as $\Pi$. Let $\Pi^{\perp}$ be the totally geodesic plane containing $\alpha_{x_{0}}(t_{0})$ and the shortest geodesic segments $\tau$ and $\tau^{\prime}$ joining it to each of $\Pi$ and $\Pi^{\prime}$, respectively. Then $\Pi^{\perp}$ intersects $\Pi$ and $\Pi^{\prime}$ orthogonally in geodesics $\gamma$ and $\gamma^{\prime}$, respectively, and there is a pentagon $P$ in $\Pi^{\perp}$ with sides consisting of $\tau$ and $\tau^{\prime}$, segments of $\gamma$ and $\gamma^{\prime}$, and the shortest geodesic segment $\lambda$ joining $\gamma$ to $\gamma^{\prime}$.
Note that $\lambda$ is necessarily orthogonal to each of $\gamma$ and $\gamma^{\prime}$, hence also to $\Pi$ and $\Pi^{\prime}$, so it is a lift of a return path of $N$. Its length is at most $2t_{0}<2h\leq\ell_{2}$, so $\lambda$ must have length $\ell_{1}$, which must be strictly less than $\ell_{2}$ in this situation. Hence $\lambda$ is a lift of $\lambda_{1}$ and therefore is the center of a lift of the muffin $M$. The pentagon $P$ is symmetric under a reflection fixing $\alpha_{x_{0}}(t_{0})$ and exchanging the sides containing it. It is divided into two quadrilaterals by the axis of this reflection. Let $Q$ be the resulting Lambert quadrilateral that contains $x_{0}$, and let $r_{0}$ be the length of its side joining $x_{0}$ to an endpoint of $\lambda$. The sides abutting this one have lengths $\ell_{1}/2$ and $t_{0}$, so the quadrilateral rule gives
$\cosh r_{0}=\frac{\tanh t_{0}}{\tanh(\ell_{1}/2)}<\frac{\tanh A}{\tanh(\ell_{1}/2)}=\cosh R.$
The inequality above comes from the hypothesis that $t_{0}<h\leq A$, and the quantity $R$ above is the one defined in (2.2.1.1). But this is the radius of the caps $C$ and $C^{\prime}$ of $M$, so the inequality implies that $x_{0}\in C\cup C^{\prime}$, contradicting our hypothesis. The claim follows.
The claim implies that the height-$h$ open collar of the preimage in $\Pi$ of $\partial N-(C\cup C^{\prime})$ is contained in $\widetilde{N}$ and does not intersect the height-$h$ open collar of any other component of $\partial\widetilde{N}$. Therefore the universal cover induces an embedding of the height-$h$ open collar of $\partial N-(C\cup C^{\prime})$ in $N$.
We now recall from Fact 3.1.7 that $U_{1}=A+W$. Here $U_{1}$, defined in Lemma 3.1.5, is a lower bound on the distance between a lift of $\lambda_{1}$ and any component of $\partial\widetilde{N}$ not containing either of its endpoints. Since $\mathrm{Muf}_{\ell_{1}}$ is contained in the $W$-neighborhood of its center, and the collar height $h$ is less than $A$, the height-$h$ open collar of $\Pi$ in $\widetilde{N}$ thus does not intersect any lift of $M$ that has no cap on $\Pi$. This implies that the height-$h$ open collar of $\partial N-(C\cup C^{\prime})$ in $N$ is disjoint from $M$. ∎
The second main result of this section is a counterpart to Proposition 3.1.11 that gives a criterion for setting the height of a collar of $\partial N$ that interacts well with a copy of $M(\ell_{2},S)$.
Proposition 3.2.3.
Regard $R$ as in (2.2.1.1) and $S=X_{12}^{1}-R$ as in Lemma 2.2.1, for $X_{ij}^{k}$ as in (2.1.7.1), as functions of variables $x=\cosh\ell_{1}$ and $y=\cosh\ell_{2}$. Suppose $N$ is an orientable hyperbolic $3$-manifold with $\partial N$ compact, connected, totally geodesic and of genus $2$, such that the lengths $\ell_{1}$ and $\ell_{2}$ of its shortest and second-shortest return paths satisfy
$a\leq\cosh\ell_{1}\leq b\leq 2\quad\mbox{and}\quad c\leq\cosh\ell_{2}\leq d\leq 3,$
and let $M$ denote the projection to $N$ of a copy of $M(\ell_{2},S)$ in $\widetilde{N}$ centered at a lift of $\lambda_{2}$. Let $U_{2}^{0}$ and $W_{S}^{0}$ be defined as in Proposition 3.1.11, and set $H^{0}=U_{2}^{0}-W^{0}_{S}$. If $H^{0}>0$, the intersection of the interior of $M$ with an open collar of $\partial N$ in $N$ is the projection to $N$ of a union of open collars of the caps of $M(\ell_{2},S)$.
Proof.
Let $\widetilde{M}$ be a copy of $M(\ell_{2},S)$ in $\widetilde{N}$ centered at a lift $\tilde{\lambda}_{2}$ of $\lambda_{2}$, and let $\Pi_{1}$ and $\Pi_{2}$ be the components of $\partial\widetilde{N}$ containing the caps of $\widetilde{M}$. For $i=1$ or $2$, since $\widetilde{M}$ is contained in the preimage of its cap in $\Pi_{i}$ under the orthogonal projection $\mathbb{H}^{3}\to\Pi_{i}$ it intersects a height-$H^{0}_{2}$ collar neighborhood of $\Pi_{i}$ in a open collar of its cap in $\Pi_{i}$. Therefore the projection of $\widetilde{M}$ to $N$ intersected with a height $H^{0}$ collar of $\partial N$ contains the union of the collar neighborhoods of its caps in $\Pi_{1}$ and $\Pi_{2}$, and any other point in this intersection lies in the projection of the intersection of $\widetilde{M}$ with the height-$H^{0}$ collar of another component $\Pi_{3}$ of $\partial\widetilde{N}$. We will show below that the latter intersection is empty, for any given such $\Pi_{3}$.
Recall from the proof of Lemma 3.1.8 that $\widetilde{M}$ is contained in a radius-$W_{2}$ neighborhood of its center $\tilde{\lambda}_{2}$, where $W_{2}$ is its waist radius, and from Lemma 3.1.5 that the quantity $U_{2}$ defined in that lemma is a lower bound for the distance from $\tilde{\lambda}_{2}$ to $\Pi_{3}$. From the proof of Proposition 3.1.11, for $(x,y)$ in the rectangle $a,b]\times[c,d]$, $W_{2}$ is bounded above by $W^{0}_{S}$ and $U_{2}$ is bounded below by $U_{2}^{0}$. Therefore by the triangle inequality, no point lies in both $\widetilde{M}$ and a height-$H^{0}$ open collar of $\Pi_{3}$. ∎
4. Volume bounds
It is recorded in Lemma 3.3 of [27] that the volume of $\mathrm{Muf}_{\ell_{1}}$ decreases with $\ell_{1}$. The Lemma below records a related derivative for later reference.
Lemma 4.1.
For the muffin $\mathrm{Muf}_{\ell_{1}}$ from Definitions and Remarks 3.1.6, depending on a parameter $\ell_{1}$, let $\mathit{VM}(x)$ record its volume as a function of $x=\cosh\ell_{1}$. Then
(4.1.1) $\mathit{VM}^{\prime}(x)=2\pi A\frac{d}{dx}(\cosh R)<0$
where $R$ as in (2.2.1.1), defined by $\cosh R(x)=\sqrt{1+1/(2x-2)}$ as a function of $x$, is the cap radius of $\mathrm{Muf}_{\ell_{1}}$ and $A$ is its altitude length, defined by $\cosh(2A)=(4x+1)/3$.
Proof.
The formula above for $R$ comes from (2.2.1.1), the one for $A$ can be deduced from (3.1.1.1), and each matches one given on the first page of [27, §3]. According to Lemma 3.3 of [27], for $A$ and $R$ as above, we have:
(4.1.2) $\mathrm{vol}(\mathrm{Muf}_{\ell_{1}})=2\pi\left(A\cosh R-\frac{\ell_{1}}{2}\right)=\pi\left((2A)\cosh R-\cosh^{-1}(x)\right).$
(As is pointed out in the proof of [27, Lemma 3.3], this equality is included in the formula stated on p. 213 of [15], which itself records an old result in hyperbolic geometry.)
Taking a derivative with respect to $x$ we obtain:
$\mathit{VM}^{\prime}(x)=\pi\left[(2A)[\cosh R]^{\prime}+(2A)^{\prime}\cosh R-\frac{1}{\sqrt{x^{2}-1}}\right]=2\pi A[\cosh R]^{\prime}$
The second equality above uses a calculus computation showing that $(2A)^{\prime}\cosh R=\frac{1}{\sqrt{x^{2}-1}}$. ∎
Proposition 4.2.
Let $N$ be an orientable hyperbolic $3$-manifold with $\partial N$ compact, connected, totally geodesic, and of genus $2$, and let $x=\cosh\ell_{1}$ and $y=\cosh\ell_{2}$, where $\ell_{1}$ and $\ell_{2}$ are the respective lengths of the shortest and second-shortest return paths of $N$. For $x\in[1.24,1.5]$, define a piecewise-constant function $Y$ of $x$ as follows:
  • for $x\in[1.24,1.25)$, $Y(x)=1.986$;
  • for $x\in[1.25,1.27)$, $Y(x)=1.9$;
  • for $x\in[1.27,1.3)$, $Y(x)=1.8$;
  • for $x\in[1.3,1.35)$, $Y(x)=1.68$;
  • for $x\in[1.35,1.4)$, $Y(x)=1.59$;
  • for $x\in[1.4,1.45)$, $Y(x)=1.55$; and
  • for $x\in[1.45,1.5)$, $Y(x)=1.52$.
If $x\in[1.24,1.5)$ and $y\geq Y(x)$, or if $x\geq 1.5$, then $\mathrm{vol}(N)\geq 7.4$.
Proof.
To prove this we will use the fundamental volume bound of [27], a result which implicitly relied on Proposition 3.2.2, and which we recorded as Lemma 3.2 in our prior work [14]. Taking $R$ as in (2.2.1.1) as a function of $x=\cosh\ell_{1}$, this is:
(4.2.1) $\mathrm{vol}(N)\geq\mathit{VM}(x)+\pi(2-\cosh R)(2H+\sinh(2H)),$
where $\mathit{VM}(x)$ records the volume of $\mathrm{Muf}_{\ell_{1}}$ as in Lemma 4.1, and for $A$ equal to the altitude length of $\mathrm{Muf}_{\ell_{1}}$ as described there, $H=\min\{A,\ell_{2}/2\}$.
Let us regard the right-hand side of (4.2.1) as defining a function $V(x,H)$ of two independent variables. For fixed $x$, this function clearly increases with $H$. Moreover, the formula for $A$ in Lemma 4.1 defines it as an increasing function of $x$, and a computation shows that when $x=1.24$, $\cosh(2A)=1.98\bar{6}$. Therefore if $N$ is as in the Proposition’s hypotheses, with $x=\cosh\ell_{1}\in[1.24,1.5]$ and $y=\cosh(\ell_{2})\geq Y(x)$, for $Y(x)$ defined in the Proposition, then $H\geq\cosh^{-1}(Y(x))/2$ and therefore $\mathrm{vol}(N)\geq V(x,\cosh^{-1}(Y(x))/2)$.
We claim that for any $x\in[1.24,1.5]$, the values of $V(x,\cosh^{-1}(Y(x))/2)$ on any interval of $x$ where $Y(x)$ is constant are minimized at the left endpoint of that interval. This follows from a computation. The second equality below uses the conclusion of Lemma 4.1.
$\frac{\partial V}{\partial x}(x,H)=\mathit{VM}^{\prime}(x)-\pi[\cosh R]^{\prime}(2H+\sinh(2H))=\pi[\cosh R]^{\prime}\left(2A-2H-\sinh(2H)\right).$
Since $\cosh R$ decreases with $x$, this quantity is positive as long as $2A<2H+\sinh(2H)$. For $x\leq 1.5$ we have $2A<\cosh^{-1}(2.34)<1.5$ and for $H\geq\cosh^{-1}(1.5)/2$,
$2H+\sinh(2H)\geq\cosh^{-1}(1.5)+\sqrt{1.5^{2}-1}>2.08.$
Therefore since $A=A(x)$ is increasing and $Y(x)$ is nonincreasing on the interval $[1.24,1.5]$, $\partial V/\partial x(x,Y(x))>0$ for all $x$ here, and the claim follows.
$x=\cosh\ell_{1}$ $Y(x)$ $V(x,H)$
$1.24$ $1.986$ $7.406$
$1.25$ $1.9$ $7.406$
$1.27$ $1.8$ $7.438$
$1.3$ $1.68$ $7.407$
$1.35$ $1.59$ $7.407$
$1.4$ $1.55$ $7.433$
$1.45$ $1.52$ $7.431$
Table 1. Values of $V(x,H)$, truncated after 33 decimal places, for $H=\cosh^{-1}(Y(x))/2$.
Table 1 collects the values $V(x,H)$, truncated after the third decimal place, for $H=\cosh^{-1}(Y(x))/2$ at left endpoints of the intervals where $Y(x)$ is constant. Since these values are all greater than $7.4$, by the claim and the observations above we have proven the Proposition’s assertions when $x=\cosh\ell_{1}\in[1.24,1.5)$.
It remains to address the case $x\geq 1.5$. Here we recall that by definition, $\ell_{2}\geq\ell_{1}$, and since $\ell_{1}/2<A$, it follows from (4.2.1) that $\mathrm{vol}(N)\geq V(x,\ell_{1}/2)$. This is exactly the function called $V$ in the proof of [14, Proposition 3.7] where it is shown to be increasing for $\cosh\ell_{1}\geq 1.439$. As its value at $\cosh\ell_{1}=1.5$ is $7.429$, truncated after three decimal places, we have that $\mathrm{vol}(N)\geq 7.4$ whenever $\cosh\ell_{1}\geq 1.5$.∎
We lack a volume bound for a manifold $N$ with totally geodesic boundary of genus $2$ if $x=\cosh\ell_{1}<1.24$ or if $x\in[1.24,1.5)$ and $\max\{\cosh E,\cosh M,\cosh\ell_{1}\}\leq\cosh\ell_{2}\leq Y(x)$, where $Y$ is defined as above and $E$ and $M$ are defined as in Proposition 2.2.2. The graphs of $\cosh E$, $\cosh M$, $\cosh\ell_{1}$ and $Y(x)$ as functions of $x=\cosh\ell_{1}$ are pictured in Figure 4.1.
[…]
Figure 4.1. $Y(x)$ versus $\max\{\cosh E,\cosh M,\cosh\ell_{1}\}$
We address these regions with double-muffin volume bounds. First we record bounds on the volume of $M(\ell_{2},S)$ over a rectangle of possible $\ell_{1}$- and $\ell_{2}$-values.
Lemma 4.3.
Regard $R$ as in (2.2.1.1) and $S=X_{12}^{1}-R$ as in Lemma 2.2.1, for $X_{ij}^{k}$ as in (2.1.7.1), all as functions of $x=\cosh\ell_{1}$ and $y=\cosh\ell_{2}$. If $\ell_{1}$ and $\ell_{2}$ satisfy
$a\leq\cosh\ell_{1}\leq b\leq 2\quad\mbox{and}\quad c\leq\cosh\ell_{2}\leq d\leq 3,$
then
(4.3.1) $\mathrm{vol}(M(\ell_{2},S))\geq\mathit{VM}_{2}^{0}\doteq 2\pi(A^{0}\cosh S(a,d)-\cosh^{-1}(d)/2),$
where $A^{0}=\tanh^{-1}\left(\cosh\left(S(a,d)\right)\sqrt{\frac{c-1}{c+1}}\right)$.
Proof.
$M(\ell_{2},S)$ has side altitude of length $A=\tanh^{-1}(\cosh S\tanh(\ell_{2}/2))$ by the formula (3.1.1.1). By Lemma 3.1.10, as a function $x$ and $y$, $S$ increases with $x$ and decreases with $y$ on the region in question. Hence it is bounded below by $S(a,d)$ there. It follows that $A$ is bounded below by $A^{0}$ as defined above on this region.
The volume of $M(\ell_{2},S)$ is given by the formula stated on p. 213 of [15], which we have already quoted in connection with (4.1.2). In the present context, the formula gives
(4.3.2) $\mathrm{vol}(M(\ell_{2},S))=2\pi(A\cosh S-\ell_{2}/2),$
By combining (4.3.2) with the lower bounds on $A$ and $S$ established above, we now obtain the lower bound on $\mathrm{vol}(M(\ell_{2},S))$ given in the Lemma’s statement. ∎
We now combine the volume of $M(\ell_{2},S)$ with existing ingredients to give the main double-muffin volume bound:
Proposition 4.4.
Suppose $N$ is an orientable hyperbolic $3$-manifold with $\partial N$ compact, connected, totally geodesic and of genus $2$, such that the lengths $\ell_{1}$ and $\ell_{2}$ of its shortest and second-shortest return paths satisfy
$a\leq\cosh\ell_{1}\leq b\leq 2\quad\mbox{and}\quad c\leq\cosh\ell_{2}\leq d\leq 3.$
For $W_{R}^{0}$, $W^{0}_{S}$, $T_{12}^{0}$ and $T_{22}^{0}$ as in (3.1.11), if $W^{0}_{R}+W^{0}_{S}\leq T^{0}_{12}$ and $2W_{S}^{0}\leq T^{0}_{22}$, then $\mathrm{vol}(N)\geq\mathit{VM}(b)+\mathit{VM}_{2}^{0}$, where $\mathit{VM}(x)$ is as in Lemma 4.1 and $\mathit{VM}_{2}^{0}$ is defined in (4.3.1).
Moreover, for $H^{0}$ as defined in Proposition 3.2.3 and $A(x)$ as in Lemma 4.1, taking $H=\min\{A(a),\cosh^{-1}(c)/2,H^{0}\}$ we have:
(4.4.1) $\mathrm{vol}(N)\geq\mathit{VM}(b)+\mathit{VM}_{2}^{0}+\pi(3-\cosh R(a)-\cosh S(b,c))\left(2H+\sinh(2H)\right),$
where $R=R(x)$ from (2.2.1.1) and $S=X_{12}^{1}-R$ from Lemma 2.2.1.
Proof.
Given the inequalities relating $W_{R}^{0}$, $W^{0}_{S}$, $T_{12}^{0}$ and $T_{22}^{0}$, Proposition 3.1.11 implies that the the universal cover $\widetilde{N}\to N$ embeds the interior of a copy of $M(\ell_{2},S)$, centered at a lift of the second-shortest return path $\lambda_{2}$, disjointly from the interior of the copy $\mathrm{Muf}_{\ell_{1}}$ centered at $\lambda_{1}$. Therefore the volume of $N$ is at least the sum of these two muffins’ volumes.
Recall from Lemma 4.1 that the function $\mathit{VM}(x)$, which records the volume of $\mathrm{Muf}_{\ell_{1}}$ as a function of $x=\cosh\ell_{1}$, is decreasing. Thus since $x\leq b$, $\mathit{VM}(x)\geq\mathit{VM}(b)$. By Lemma 4.3, $\mathit{VM}_{2}^{0}$ bounds the volume of $M(\ell_{2},S)$ below for these values of $\ell_{1}$ and $\ell_{2}$. Thus $\mathrm{vol}(N)\geq\mathit{VM}(b)+\mathit{VM}_{2}^{0}$.
We can augment this lower bound by adding the volume of a collar of the region in $\partial N$ outside the union of the caps of the embedded copies of $\mathrm{Muf}_{\ell_{1}}$ and $M(\ell_{2},S)$, as long as the collar is not too high. If $B$ is the area of the region in $\partial N$ and $H$ is the collar height, then the collar volume $V$ satisfies $V=B\cdot(2H+\sinh(2H))/4$. As $\partial N$ has area $4\pi$, by the Gauss-Bonnet theorem, and a hyperbolic disk of radius $r$ has area $2\pi(\cosh r-1)$, we have:
$B=4\pi-4\pi(\cosh R-1)-4\pi(\cosh S-1)=4\pi\left(3-\cosh R-\cosh S\right).$
A collar height of $H=\min\{A,\ell_{2}/2\}$ was used in [27] and [14], where $A=A(x)$ is defined by $\cosh(2A)=(4x+1)/3$ as in Lemma 4.1. For a copy of $\mathrm{Muf}_{\ell_{1}}$ in $N$ centered at $\lambda_{1}$, with caps $U$ and $U^{\prime}$, Proposition 3.2.2 asserts that a height-$H$ open collar of $\partial N-(U\cup U^{\prime})$ is is embedded in $N$ disjointly from $\mathrm{Muf}_{\ell_{1}}$. Since $A$ is an increasing function of $x$, the quantity $\min\{A(a),\cosh^{-1}(c)/2\}$ bounds $\min\{A,\ell_{2}/2\}$ below on the entire rectangle.
For $H^{0}$ as defined in Proposition 3.2.3, that result implies that a height-$H^{0}$ collar of $\partial N$ intersects the embedded copy of $M(\ell_{2},S)$ in the union of open collars of its caps $V$ and $V^{\prime}$ (noting that the bounds on $\ell_{1}$ and $\ell_{2}$ in the hypothesis of the present proposition match those in the hypothesis of Proposition 3.2.3). Therefore a collar of $\partial N-(U\cup U^{\prime}\cup V\cup V^{\prime})$ with height $H=\min\{A(a),\cosh^{-1}(c)/2,H^{0}\}$ is embedded in $N$ without overlapping either the copy of $\mathrm{Muf}_{\ell_{1}}$ or of $M(\ell_{2},S)$, so it contributes volume
$V=\pi(3-\cosh R-\cosh S)(2H+\sinh(2H))$
to the volume of $N$, independently of the muffins.
In order to bound $V$ below in terms of the given bounds on $x=\cosh\ell_{1}$ and $y=\cosh\ell_{2}$, we bound $R$ and $S$ above. Since $R$ decreases with $x$, its value is bounded above by $R(a)$. And by Lemma 3.1.10, the value of $S$ is bounded above by $S(b,c)$. Together with the height bound $H$ described above, this gives the collar volume’s contribution to the Proposition’s lower bound on $\mathrm{vol}(N)$. ∎
We use Proposition 4.4 and a numerical scheme, whose output is summarized in the picture below, to complete the proof of the first main result of this section.
[…]
Figure 4.2. Some volume bounds
Theorem 4.5.
Let $N$ be an orientable hyperbolic $3$-manifold with $\partial N$ compact, connected, totally geodesic, and of genus $2$. If the length $\ell_{1}$ of the shortest return path of $N$ satisfies $\cosh\ell_{1}\geq 1.24$ then $N$ has volume at least $7.4$.
Proof.
Given Proposition 4.2, to prove the Theorem we must bound the volume of a manifold $N$ whose shortest return path length $\ell_{1}$ satisfies $1.24\leq\cosh\ell_{1}\leq 1.5$, and whose second-shortest return path length $\ell_{2}$ satisfies $\cosh\ell_{2}\leq Y(x)$, for the piecewise-constant function $Y$ of the Proposition. For this we implement the volume bounds of Proposition 4.4 numerically as follows.
On each maximal subinterval $a_{0},b_{0})$ of $[1.24,1.5)$ where $Y$ is constant, we find the minimum value of the function $\max\{\ell_{1},E,M\}$ from Lemma 2.2.2. Let $c_{0}$ be the hyperbolic cosine of this value, truncated after three decimal places. By Lemma 2.2.2, for $N$ with $\cosh\ell_{1}\in[a_{0},b_{0}]$, $\cosh\ell_{2}\geq c_{0}$. Let $d_{0}$ be the value of $Y$ on $a_{0},b_{0})$. We have written Python scripts that divide the rectangle $a_{0},b_{0}]\times[c_{0},d_{0}]$ into subrectangles of size $0.001\times 0.001$, and on each subrectangle $a,b]\times[c,d]$ performs the following tasks.
  1. (1)
    Check whether $a,b]\times[c,d]$ intersects the possible subregion: the subset of $a_{0},b_{0}]\times[c_{0},d_{0}]$ consisting of $(x,y)$-values that satisfy the criterion of Lemma 2.2.2. For instance, $E(x)=\max\{\ell_{1},E,M\}$ for $x\leq 1.366$, so since $E$ is a decreasing function of $x$ here (cf. [14, Lemma 3.4]), for $b\leq 1.366$ this is tantamount to checking that $d\geq\cosh E(b)$.
    If $a,b]\times[c,d]$ does not intersect the possible subregion then we ignore it since in this case, by Lemma 2.2.2, there does not exist a manifold satisfying the hypotheses of the present result and with an $(x,y)$-value in $a,b]\times[c,d]$.
  2. (2)
    If $a,b]\times[c,d]$ does intersect the possible subregion, check the muffin-embedding criteria of Proposition 3.1.11. These hold in each such case, confirming the expectation recorded above that result.
  3. (3)
    On $a,b]\times[c,d]$ intersecting the possible subregion, compute the two volume bounds supplied by Proposition 4.4, recording the larger. (The reason that the sum of muffin volumes $\mathit{VM}(b)+\mathit{VM}^{0}_{2}$ may be larger than the bound (4.4.1), which also incorporates the volume of a collar of a region in $\partial N$, is that in some cases the given lower bound $3-\cosh R(a)-\cosh S(b,c)$ for the area of that region may be negative.)
The Python scripts performing this task are included in the ancillary files as VolScript_E.py, VolScript_M.py, and VolScript_x.py, where the different names refer to the different lower bounds that the scripts call for $\max\{\ell_{1},E,M\}$ (each choice being best for a particular range of $x$ values). They call functions from the script formulas.py, which collects relevant formulas from elsewhere in this paper. Their output is summarized in text files, also included in the ancillary materials, such as 124_125.txt: this one recording bounds for possible $(x,y)$-values with $x\in[1.24,1.25]$.
The resulting volume bounds are summarized in Figure 4.2. Each big rectangle $a_{0},b_{0}]\times[c_{0},d_{0}]$ from above is shaded in the Figure, and the smallest of the volume bounds produced by the process above—taken over all subrectangles $a,b]\times[c,d]$—is recorded in or directly above the big rectangle. (The subrectangle on which the minimum is attained is also blackened in the Figure.) Since each of these bounds is larger than $7.4$, the Theorem follows from Proposition 4.2. ∎
The next theorem provides a volume bound for manifolds satisfying an extra condition.
Theorem 4.6.
Let $N$ be an orientable hyperbolic $3$-manifold with $\partial N$ compact, connected, totally geodesic, and of genus $2$. If the length $\ell_{1}$ of the shortest return path of $N$ satisfies $\cosh\ell_{1}\leq 1.24$ and $\widetilde{N}$ contains no $(1,1,1)$-hexagon then $N$ has volume at least $7.409$.
Proof.
Since $\widetilde{N}$ contains no $(1,1,1)$-hexagon, Proposition 2.3.1 asserts that $x\doteq\cosh\ell_{1}>1.23$. The proof will follow the same scheme as that of Theorem 4.5, but we will use the interval $[1.23,1.24]$ of possible values of $x$, and will appeal to Proposition 2.3.1 again to get a non-constant upper bound $Y_{0}(x)$ on $\cosh\ell_{2}$ in terms of $x$. From the formula of Proposition 2.3.1 we obtain $Y_{0}(x)=(2+2\sqrt{3})x^{2}-(3+2\sqrt{3})$.
We use a single large rectangle $1.23,1.24]\times[1.74,1.938]$. The lower bound of $y=1.74$ is the minimum of $\cosh E(x)$ on this interval, which is attained at $x=1.24$, truncated after three decimal places. The upper bound is the maximum of $Y_{0}(x)$ on the interval, also attained at $1.24$, rounded up after three decimal places. Because $Y_{0}(x)$ is not constant here but increasing, we alter step (1) above for a subrectangle $a,b]\times[c,d]$ by checking that both $d\geq\cosh E(b)$ and $c\leq Y_{0}(x)$.
To achieve a slight improvement in our volume bounds, we use a subrectangle size of $0.0005\times 0.001$—that is, we divide the previous subrectangles in half vertically. The smallest bound resulting from Theorem 4.5’s revised process is $7.409$, truncated after three decimal places, attained on the subrectangle $1.2305,1.231]\times[1.808,1.809]$, proving the result. (Using the previous subrectangle size would yield $7.399$ as a bound.) The relevant Python script is VolScript_EUl2.py, and its output is 123_124.txt. ∎
Theorems 4.5 and 4.6 together yield an unconditional lower bound on volume in the absence of a $(1,1,1)$-hexagon.
Corollary 4.7.
Let $N$ be an orientable hyperbolic $3$-manifold with $\partial N$ compact, connected, totally geodesic, and of genus $2$. If the universal cover $\widetilde{N}$ of $N$ contains no $(1,1,1)$-hexagon then $\mathrm{vol}(N)\geq 7.4$.
Proof.
For a manifold $N$ satisfying the hypotheses of the corollary, if the length $\ell_{1}$ of the shortest return path of $N$ satisfies $\cosh\ell_{1}\leq 1.24$ then $\mathrm{vol}(N)\geq 7.409$ by Theorem 4.6. If $\cosh\ell_{1}\geq 1.24$ then $\mathrm{vol}(N)\geq 7.4$ by Theorem 4.5. Thus $\mathrm{vol}(N)$ is at least $7.4$ in both cases.∎
5. Trimonic Manifolds from a $(1,1,1)$-hexagon
Here we strengthen some results from Section 6 of [14]. Our first result removes a restriction on the length of $\ell_{1}$ from the hypothesis of Lemma 6.6 there.
Lemma 5.1.
Let $N$ be a hyperbolic $3$-manifold with compact totally geodesic boundary. If $C$ is a $(1,1,1)$-hexagon in $\widetilde{N}$ and $\tilde{\lambda}_{1}$ is a lift of the shortest return path of $N$, then $\tilde{\lambda}_{1}$ is an internal edge of $C$ or $\tilde{\lambda}_{1}\cap C=\emptyset$.
We recall from Definition 2.1.5 here that the three internal edges of a $(1,1,1)$-hexagon are short cuts, i.e. lifts of the shortest return path, and its other, external edges lie in components of $\partial\widetilde{N}$. This matches the usage in [14].
Proof.
We follow the proof of Lemma 6.6 of [14]: let $\Pi_{1}$ and $\Pi_{2}$ be the components of $\partial\widetilde{N}$ containing the endpoints of $\tilde{\lambda}_{1}$ and $\Pi$ the geodesic plane containing $C$. As shown in the proof of [14, L. 6.6]’s first paragraph, if $\tilde{\lambda}_{1}$ lies in $\Pi$ then it is an internal edge of $C$. (This argument does not require a condition on the length of $\tilde{\lambda}_{1}$.) We therefore suppose below that $\tilde{\lambda}_{1}$ does not lie in $\Pi$ and hence intersects it transversely in a point $x$.
As in the previous proof, we take $\Pi^{\prime}$ to be the component of $\partial\widetilde{N}$ containing the external edge of $C$ closest to $x$. As observed there, the distance from $\Pi^{\prime}$ to $x$ is at most $A$, defined in equation (3.1.1) of [14] by $\cosh A=\sqrt{\frac{2}{3}(\cosh\ell_{1}+1)}$ (the same “AA” from Lemma 4.1 here). This thus bounds the distance from $\Pi^{\prime}$ to $\tilde{\lambda}_{1}$ above.
On the other hand, we repeat for emphasis that $\Pi^{\prime}\neq\Pi_{i}$ for $i=1$ or $2$ (this was also used in [14, L. 6.6]), since $\tilde{\lambda}_{1}$ intersects $C$ transversely but intersects $\Pi_{1}$ and $\Pi_{2}$ at right angles. Therefore the distance from $\Pi^{\prime}$ to $\tilde{\lambda}_{1}$ is bounded below by the quantity $U_{1}$ from Lemma 3.1.5 of this paper, which is given strictly in terms of $\cosh\ell_{1}$ by:
$\cosh U_{1}=\sqrt{\frac{2\cosh^{2}\ell_{1}}{\cosh\ell_{1}-1}+1}=\sqrt{\frac{2\cosh\ell_{1}-1}{\cosh\ell_{1}-1}(\cosh\ell_{1}+1)}$
A little manipulation then shows that $\cosh U_{1}>\sqrt{2(\cosh\ell_{1}+1)}$. It follows that $U_{1}>A$, yielding a contradiction. Thus $\tilde{\lambda}_{1}$ can only intersect $C$ as one of its internal edges.∎
Lemma 5.1 in turn facilitates an analogous strengthening of [14, Lemma 6.7]:
Lemma 5.2.
Let $N$ be a hyperbolic $3$-manifold with compact totally geodesic boundary. If $C$ and $C^{\prime}$ are distinct $(1,1,1)$-hexagons in $\widetilde{N}$, then $C\cap C^{\prime}$ is empty or a single internal edge of each.
Proof.
The proof of [14, Lemma 6.7] still goes through here (i.e. without the prior result’s hypothesis that $\cosh\ell_{1}\leq 1.215$): one need only replace the reference to [14, Lemma 6.6] in the proof of [14, Lemma 6.7] by a reference to Lemma 5.1 above. ∎
The following result strengthens both Propositions 6.8 and 6.9 of [14] by considerably weakening the restriction on $\ell_{1}$. The term trimonic manifold, which appears in the following statement, is defined in [14, Definition 5.7], while the term non-degenerate is defined in [14, Definition 5.10].
Proposition 5.3.
Let $N$ be an orientable hyperbolic $3$-manifold with $\partial N$ compact, connected, totally geodesic, and of genus $2$, such that there is a $(1,1,1)$-hexagon in $\widetilde{N}$ and
$\cosh\ell_{1}<\frac{\cos(2\pi/9)}{2\cos(2\pi/9)-1}=1.43969$...
Then there is a submanifold $X\subset N$ with $\partial N\subset X$, such that $X$ is a non-degenerate trimonic manifold relative to $\partial N$.
Proof.
The trimonic manifold $X$ is constructed in the proof of [14, Prop. 6.8] as a regular neighborhood of the union of $\partial N$ with the image in $N$ of a $(1,1,1)$-hexagon $C$ in $\widetilde{N}$ under the universal covering projection $\widetilde{N}\to N$. Following that proof, we denote by $f$ the restriction of the covering projection to $C$. The upper bound on $\cosh\ell_{1}$ ensures that $N$ has a unique shortest return path $\lambda_{1}$, by [14, Lemma 6.3], and hence that $f$ projects every internal edge of $C$ to $\lambda_{1}$.
In order to show that $X$ is a trimonic manifold relative to $\partial N$, we must verify that it has the properties (1)–(6) stated in [14, Dfn. 5.7]. Property (1), which in the proof of [14, Prop. 6.8] was verified by an appeal to Lemma 6.6 of [14], follows here from Lemma 5.1 above. Lemma 6.7 of [14], which in the earlier argument was used to establish Property (3), is replaced here by Lemma 5.2. The verifications of Properties (2), (4), (5) and (6) go through exactly as in the earlier argument. Furthermore, the argument for non-degeneracy given in Proposition 6.9 of [14] does not depend on any particular upper bound for $\cosh\ell_{1}$ and so carries through without alteration. ∎
6. Background on topology and least-area surfaces
The proof of Theorem 7.2, which was stated in the Introduction, combines the results of the preceding sections with concepts and results from three-manifold topology and geometry which will also be important in later sections. We introduce these here. Subsection 6.1 establishes topological conventions, definitions and observations that will be used throughout the rest of the paper. Subsection 6.2 reviews a series of deep differential geometric results on minimal surfaces in three-manifolds proved by other authors, which were used in the proof of [14, Th. 7.4], and adapts them to the present context. Their first application here is in the proof of Theorem 7.1, which is itself an ingredient of the proof of Theorem 7.2.
6.1. Topological background
The material concerning manifolds in this subsection, and elsewhere in the paper, is to be understood in the smooth category. However, we shall often implicitly exploit the equivalence of the smooth and piecewise-linear categories in dimension $3$ to go back and forth between the two categories.
We stipulate, as part of the definition of connectedness, that a connected topological space is non-empty.
The Euler characteristic of a finitely triangulable space $Y$ will be denoted $\chi(Y)$, and we will set $\bar{\chi}(Y)=-\chi(Y)$.
If $A$ is a subset of a topological space $X$, we denote the frontier of $A$ in $X$, defined to be $\overline{A}\cap\overline{X-A}$, by $\mathop{\rm Fr}_{X}A$.
We shall follow the conventions of [32] regarding Heegaard splittings. In particular, each compact, connected, orientable $3$-manifold-with-boundary $M$ has a well-defined Heegaard genus, which we denote by ${\rm Hg}(M)$.
A connected submanifold $Y$ of a connected manifold $X$ is said to be $\pi_{1}$-injective in $X$ if the inclusion homomorphism $\pi_{1}(Y)\to\pi_{1}(X)$ is injective. More generally, a (possibly empty) submanifold $Y$ of a manifold $X$ is said to be $\pi_{1}$-injective in $X$ if each component $C$ of $Y$ is $\pi_{1}$-injective in the component of $X$ containing $C$.
A $3$-manifold $M$ is said to be irreducible if $M$ is connected and every (tame) $2$-sphere in $M$ bounds a ball. One says that a two-dimensional submanifold $\mathcal{S}$ of an irreducible $3$-manifold $M$ is incompressible if (1) $S$ is closed and orientable, and is contained in $\mathop{\rm int}M$, (2) no component of $S$ is a $2$-sphere, and (3) $\mathcal{S}$ is $\pi_{1}$-injective. The $3$-manifold $M$ is said to be boundary-irreducible if $\partial M$ is $\pi_{1}$-injective, and boundary-reducible otherwise.
In this paper, we say that a connected $3$-manifold is strongly atoral if its fundamental group has no rank-$2$ free abelian subgroup. We shall say that a $3$-manifold is simple if it is compact, orientable, irreducible, boundary-irreducible, strongly atoral, and has an infinite fundamental group. Note that a simple $3$-manifold $M$ is not homeomorphic to a ball; since $M$ is irreducible, it follows that no boundary component of $M$ can be a $2$-sphere.
Note also that the definition of “simple” given here is similar but not identical to the definition of the same term in [8]. It is easy to see that a $3$-manifold is simple in the sense defined in [8] if and only if it is either simple in the sense defined here or is homeomorphic to a ball.
Note that every closed, orientable, hyperbolic $3$-manifold is simple, and that in a simple $3$-manifold, every connected incompressible surface has genus at least $2$.
[…]
If $M$ is simple, and $S$ is incompressible and non-empty, then each component of $M\setminus\setminus S$ is simple and has non-empty boundary.
Any simple $3$-manifold $U$ with non-empty boundary has a well-defined relative characteristic submanifold $\Sigma_{U}$ in the sense of [25] and [24]. (In the notation of [24], $(\Sigma_{U},\Sigma_{U}\cap\partial U)$ is the characteristic pair of $(U,\partial U)$. The assumptions that $U$ is simple and has non-empty boundary are enough to guarantee that $(U,\partial U)$ is a “sufficiently large pair” in the sense of [24], so that the Characteristic Pair Theorem [24, p. 118] guarantees that the characteristic pair is well defined. The arguments of [24], and the corresponding arguments in [25], are done in the piecewise-linear category; as we are translating the results to the smooth category, we must regard $\Sigma_{U}$ as a smooth manifold with corners —we refer to [28, Ch. 16] for definitions and basic facts— such that $\Sigma_{U}\cap\partial U$ and $\mathop{\rm Fr}_{U}\Sigma_{U}$ are smooth manifolds with boundary.) For each component $C$ of $\Sigma_{Q}$ either (i) $C$ may be given the structure of an $I$-bundle over a compact (smooth) $2$-manifold-with-boundary $F_{C}$ with $\chi(F_{C})<0$, in such a way that $\mathop{\rm Fr}_{Q}C$ is the preimage of $\partial F_{C}$ under the bundle projection, or (ii) $C$ is homeomorphic to a solid torus and the components of $\mathop{\rm Fr}_{Q}C$ are (smooth) annuli in $\partial C$ that are homotopically non-trivial in $C$.
Now let $Q$ be a compact $3$-manifold, each of whose components is a simple manifold with non-empty boundary. We denote by $\Sigma_{Q}\subset Q$ the union of the submanifolds $\Sigma_{U}$, where $U$ ranges over the components of $Q$. Since $\Sigma_{Q}$ is a manifold with corners whose frontier is a smooth manifold with boundary, $\overline{Q-\Sigma_{Q}}$ is a manifold with corners. Each component of $\overline{Q-\Sigma_{Q}}$ either has (strictly) negative Euler characteristic, or may be identified (by a diffeomorphism of manifolds with corners) with $S^{1}\times[0,1]\times[0,1]$ in such a way that its frontier in $Q$ is $S^{1}\times[0,1]\times\{0,1\}$. (To rule out components of $\overline{Q-\Sigma_{Q}}$ with strictly positive Euler characteristic, we must show that no boundary component $W$ of $\overline{Q-\Sigma_{Q}}$ is a $2$-sphere. If $W\cap\Sigma_{Q}=\emptyset$ this follows from the observation, made above, that no boundary component of a simple $3$-manifold $Q$ is a $2$-sphere. If $W\cap\Sigma_{Q}\neq\emptyset$ then $W$ contains a component of $\mathop{\rm Fr}_{Q}\Sigma_{Q}$, which is a $\pi_{1}$-injective annulus in $Q$ and therefore cannot be contained in a $2$-sphere.) We denote by $\mathop{\rm kish}(Q)$ the union of all components of $\overline{Q-\Sigma_{Q}}$ that have negative Euler characteristic, and set $\mathop{\rm kish}^{0}(Q)=\mathop{\rm kish}Q-\mathop{\rm Fr}_{Q}\mathop{\rm kish}Q$. Thus $\mathop{\rm kish}Q$ is a $3$-manifold-with-corners, while $\mathop{\rm kish}^{0}Q$ is a smooth $3$-manifold-with-boundary; and, by definition, for each component $K$ of $\mathop{\rm kish}(Q)$, we have $\chi(K)<0$, or equivalently $\bar{\chi}(K)\geq 1$.
To say that $Q$ is acylindrical (where $Q$ is still understood to be a compact $3$-manifold, each of whose components is a simple manifold with non-empty boundary) means that $\Sigma_{Q}=\emptyset$; this is equivalent to saying that $\mathop{\rm kish}(Q)=Q$.
Note that every compact hyperbolic $3$-manifold with non-empty totally geodesic boundary is (i) simple and (ii) acylindrical.
As in [3], we define a book of $I$-bundles to be a triple ${\mathcal{W}}=(W,{\mathcal{B}},{\mathcal{P}})$, where (1) $W$ is a compact, orientable smooth $3$-manifold with boundary; (2) $\mathcal{B}$} and $\mathcal{P}$} are manifolds with corners such that ${\mathcal{B}}\cup{\mathcal{P}}=W$ and ${\mathcal{B}}\cap{\mathcal{P}}=\mathop{\rm Fr}_{W}{\mathcal{B}}=\mathop{\rm Fr}_{W}{\mathcal{P}}$, and each of the sets ${\mathcal{B}}\cap{\mathcal{P}}$, ${\mathcal{B}}\cap\partial W$, and ${\mathcal{P}}\cap\partial W$ is a smooth manifold with boundary; (3) each component $B$ of $\mathcal{B}$} is a solid torus whose frontier components in $W$ are (smooth) annuli in $\partial B$ that are homotopically non-trivial in $B$; and (4) each component $P$ of $\mathcal{P}$} may be given the structure of an $I$-bundle over a compact smooth $2$-manifold with boundary $F_{P}$ in such a way that $\mathop{\rm Fr}_{Q}P$ is the preimage of $\partial F_{P}$ under the bundle projection. The components of $\mathcal{B}$} and $\mathcal{P}$} are called bindings and pages of $\mathcal{W}$}, respectively. The manifold $W$ is called the underlying manifold of $\mathcal{W}$} and may be denoted $|{\mathcal{W}}|$. We will say that a book of $I$-bundles $\mathcal{W}$} is connected if $|{\mathcal{W}}|$ is connected.
It follows from the discussion above that if $Q$ is a compact $3$-manifold, each of whose components is a simple manifold with non-empty boundary, and if $\mathop{\rm kish}(Q)=\emptyset$, then $Q$ is the underlying manifold of some book of $I$-bundles $\mathcal{W}$}. Indeed, we may take ${\mathcal{W}}=(Q,{\mathcal{B}},{\mathcal{P}})$, where $\mathcal{B}$} is a regular neighborhood of the union of all solid torus components of $\Sigma_{Q}$, and $\mathcal{P}$} is a regular neighborhood of the union of all components of $\Sigma_{Q}$ having negative Euler characteristic.
6.2. Background on least-area surfaces
Here we review a series of deep differential geometric results on minimal surfaces in three-manifolds proved by other authors, which were used in the proof of [14, Th. 7.4]. We further adapt these results here for use in the proof of Theorem 7.1 of the present paper—which is an upgrade of [14, Th. 7.4]—and for further applications later in this paper.
Definitions and Remarks 6.2.1.
If $S$ is an oriented smooth manifold and $M$ is a Riemannian manifold, possibly with boundary, with $\dim M\geq\dim S$, any smooth immersion $f\colon\thinspace S\to M$ pulls the Riemannian metric $g$ on $M$ back to a Riemannian metric $f^{*}g$ on $S$. For any (measurable) $A\subset S$, we define the volume of $A$ under $f$, denoted by $\mathop{\rm vol}_{f}(A)$, to be the integral of the volume form of $f^{*}g$ over $A$. This extends naturally to a definition of $\mathop{\rm vol}_{f}(A)$ for any smooth map $f:S\to M$, by taking the volume form to be 0 at points where the derivative of $f$ is singular. Thus if $f$ is not an immersion, we may have $\mathop{\rm vol}_{f}(A)=0$ even if (say) $A$ has non-empty interior.
For $\dim S=2$ we will use the term “area” in place of “volume,” and write $\mathop{\rm area}_{f}(A)$ in place of $\mathop{\rm vol}_{f}(A)$ for measurable $A\subset S$. We will say that a smooth map $f\colon\thinspace S\to M$ is least-area if $\mathop{\rm area}_{f}(S)\leq\mathop{\rm area}_{g}(S)$ for any map $g\colon\thinspace S\to M$ smoothly homotopic to $f$.
There is an equivalent definition of the volume of $A\subset S$ under the smooth map $f$ in the special case where $\dim M=\dim S$ and where $M$, as well as $S$, is equipped with an orientation. In this situation, the metric on $M$ defines a volume form $\alpha$, and the pulled back form $f^{*}(\alpha)$ may be written as $h\cdot\omega$, where $\omega$ is a non-vanishing form that determines the given orientation of $S$. We then have $\mathop{\rm vol}_{f}(A)=\int_{A}|h|\omega$. In particular, since $\left|\int_{A}|h|\omega\right|\geq|\int_{A}h\omega|$, it follows that $\mathop{\rm vol}_{f}(A)\geq|\int_{A}f^{*}(\alpha)|$.
There is a still more general notion of volume of submanifolds from geometric measure theory—the Hausdorff measure—that applies even to non-smooth maps and gives the same result as the definition above for smooth embeddings.
Notation and Remarks 6.2.2.
By definition, a riemannian metric on a manifold $M$ gives an inner product, and hence a norm, on the tangent space at any point of $M$. If $\kappa:M_{1}\to M_{2}$ is a smooth map between riemannian manifolds, then for every $\mathbf{x}\in M_{1}$ the derivative $d\kappa_{\mathbf{x}}:T_{\mathbf{x}}M_{1}\to T_{\kappa(\mathbf{x})}M_{2}$ is a linear map, and has an operator norm $\|d\kappa_{\mathbf{x}}\|=\max_{v\in T_{\mathbf{x}}M_{1},\|x\|=1}\|d\kappa_{x}(v)\|$.
Now suppose that $\|d\kappa_{\mathbf{x}}\|\leq 1$ for every $\mathbf{x}\in M_{1}$. Then for any oriented smooth manifold $S$, any smooth map $f:S\to M_{1}$, and any point $\mathbf{u}\in S$ such that $df_{\mathbf{u}}$ is non-singular, the norms defined on $T_{\mathbf{u}}$ by the pullbacks via $f$ and $\kappa\circ f$ of the metrics $g_{1}$ and $g_{2}$ on $M_{1}$ and $M_{2}$ satisfy $\|(\kappa\circ f)^{*}(g_{2})\|\leq\|f^{*}(g_{2})\|$. This implies in particular that if $\beta_{1}$ and $\beta_{2}$ denote the volume forms of the respective pullbacks, and if for $i=1,2$ we choose an $n$-form $\omega$ defining the orientation of $S$ and write $\beta_{i}=h_{i}\omega$ for some function $h_{i}$ which is positive wherever $df_{\mathbf{u}}$ is non-singular, then $h_{2}\leq h_{1}$. If as in 6.2.1 we extend the $\beta_{i}$ to all of $S$ by defining them to be 0 at points where $df_{\mathbf{u}}$ is singular, and if we define the $h_{i}$ to be 0 at such points as well, then the inequality $h_{2}\leq h_{1}$ holds on all of $S$; upon integrating we conclude that for any (measurable) set $A\subset S$ we have $\mathop{\rm vol}_{\kappa\circ f}(A)\leq\mathop{\rm vol}_{f}(A)$, in the notation of 6.2.1.
If, in addition to the assumption that $\|d\kappa_{\mathbf{x}}\|\leq 1$ for every $\mathbf{x}\in M_{1}$, we assume that $\|d\kappa_{\mathbf{x}}\|<1$ for some $\mathbf{x}\in f(A)$, the same argument shows that $\mathop{\rm vol}_{\kappa\circ f}(A)<\mathop{\rm vol}_{f}(A)$.
Proposition 6.2.3.
Let $\Pi$ be any totally geodesic plane in ${\mathbb{H}}^{3}$ and let $\pi:{\mathbb{H}}^{3}\to\Pi$ denote orthogonal projection. Then if $\mathbf{x}$ is any point of ${\mathbb{H}}^{3}$, and $D$ denotes the hyperbolic distance from $\mathbf{x}$ to $\Pi$, we have $\|d\pi_{\mathbf{x}}\|=1/\cosh D$, where $d\pi_{\mathbf{x}}:T_{\mathbf{x}}{\mathbb{H}}^{3}\to T_{\pi{\mathbf{x}}}\Pi$ is the derivative map, and the operator norm $\|d\pi_{\mathbf{x}}\|$ is defined as in 6.2.2. In particular we have $\|d\pi_{\mathbf{x}}\|\leq 1$ for every $\mathbf{x}\in{\mathbb{H}}^{3}$, and $\|d\pi_{\mathbf{x}}\|<1$ for every $\mathbf{x}\in{\mathbb{H}}^{3}-\Pi$.
Proof.
Because the isometries of $\mathbb{H}^{3}$ act transitively on its collection of totally geodesic planes and conjugate orthogonal projections to orthogonal projections, we may fix a particular plane $\Pi$ on which to establish the assertions of the proposition. Using the upper half-space model for $\mathbb{H}^{3}$, we choose $\Pi$ to be the unit hemisphere centered at $\mathbf{0}$. The orthogonal projection $\pi$ to this plane is given in terms of the ambient coordinates on $\mathbb{R}^{3}$ by
$\pi(x,y,z)=\frac{1}{1+x^{2}+y^{2}+z^{2}}\,\left(2x,2y,\sqrt{(1+x^{2}+y^{2}+z^{2})^{2}-4(x^{2}+y^{2})}\right)$
Let us prove the first assertion of the proposition, that for any $\mathbf{x}\in{\mathbb{H}}^{3}$ we have $\|d\pi_{\mathbf{x}}\|=1/\cosh D$, where $D$ denotes the distance from $\mathbf{x}$ to $\Pi$.
Because the stabilizer of $\Pi$ in $\mathbb{H}^{3}$ acts transitively on the points of $\Pi$, we may assume without loss of generality that $\pi(\mathbf{x})=(0,0,1)$; after possibly modifying $\mathbf{x}$ by a reflection about $\Pi$ we may further assume that $\mathbf{x}=(0,0,z)$, where $z=e^{D}$.
Denoting the standard basis vectors for $\mathbb{R}^{3}$ as $\mathbf{e}_{1}$, $\mathbf{e}_{2}$, $\mathbf{e}_{3}$, we have that $\{z\mathbf{e}_{1},z\mathbf{e}_{2},z\mathbf{e}_{3}\}$ and $\{\mathbf{e}_{1},\mathbf{e}_{2}\}$ are orthonormal bases for $T_{\mathbf{x}}\mathbb{H}^{3}$, and $T_{\pi\mathbf{x}}\Pi$, respectively. In terms of these bases, the derivative $d\pi_{\mathbf{x}}$ of $p$ at $\mathbf{x}$ is given by
(6.2.3.1) $d\pi_{\mathbf{x}}=\frac{1}{\cosh D}\begin{pmatrix}1&0&0\\ 0&1&0\end{pmatrix}$.
The matrix in (6.2.3.1) can be obtained by computing the usual Jacobian matrix at $\mathbf{x}=(0,0,z)$, of partial derivatives of the components of $\pi$, applying it to the basis vectors for $T_{\mathbf{x}}\mathbb{H}^{3}$ given above, expressing their images in terms of the basis vectors for $T_{\pi\mathbf{x}}\Pi$, then substituting $e^{D}$ for $z$ and simplifying.
The expression (6.2.3.1) for $d\pi_{\mathbf{x}}$ immediately implies the first assertion of the proposition, that $\|d\pi_{\mathbf{x}}\|=1/\cosh D$. ∎
Corollary 6.2.4.
The inclusion map $S\hookrightarrow M$ of a totally geodesic surface $S$ in a closed hyperbolic $3$-manifold $M$ is least-area, in the sense of 6.2.1.
Proof.
Since $S$ is least-area if each of its components is, we may assume that $S$ is connected.
[…]
If $f:S\to M$ is a map homotopic to $i$, and we choose a smooth homotopy $F\colon\thinspace S\times I\to M$ from $i$ to $f$, then $F$ admits a lift $\widetilde{F}:S\to\widetilde{M}$. Now $H\doteq p\circ\pi\circ\widetilde{F}$ is a homotopy from the identity map of $S$ to $j\doteq p\circ\pi\circ\widetilde{f}:S\to S$, where $\widetilde{f}$ is a lift of $f$ to $\widetilde{M}$. Since $\|d\pi_{\mathbf{x}}\|\leq 1$ for every $\mathbf{x}\in\widetilde{M}$, it follows from 6.2.2 that $\mathop{\rm area}_{\pi\circ\widetilde{f}}(S)\leq\mathop{\rm area}_{\widetilde{f}}(S)$. Since $p$ is a local isometry, this inequality may be rewritten as $\mathop{\rm area}_{j}(S)\leq\mathop{\rm area}_{f}(S)$.
Now fix an orientation of $S$. By an observation made in 6.2.1 we have $\mathop{\rm area}_{j}(S)\geq|\int_{S}j^{*}(\alpha)|$, where $\alpha$ is the area form determined by the orientation of $S$. But it is a standard consequence of Stokes’s Theorem that the pull-backs of an $n$-form under homotopic maps between closed $n$-manifolds have the same integral. Since $j:S\to S$ is homotopic to the identity, it follows that $\int_{S}j^{*}(\alpha)=\int_{S}{\rm id}_{S}^{*}(\alpha)=\mathop{\rm area}(S)\geq 0$, and hence that $\mathop{\rm area}_{j}(S)\geq\mathop{\rm area}(S)$. Noting that since $S$ is totally geodesic, $\mathop{\rm area}_{i}(S)=\mathop{\rm area}(S)$, we conclude that $\mathop{\rm area}_{f}(S)\geq\mathop{\rm area}_{j}(S)\geq\mathop{\rm area}_{i}(S)$. ∎
The lemma below is a variation on a standard consequence of deep results of Thurston [35], Agol-Storm-Thurston [4], and Miyamoto [30]; compare it with Theorems 7.2 and 9.1 of [4].
Lemma 6.2.5.
Let $N$ be a compact $3$-manifold, each of whose components is simple and has non-empty boundary. Suppose that $N$ is equipped with a hyperbolic metric such that $\partial N$ is a minimal surface. Then we have $\mathop{\rm vol}N>V_{\rm oct}\,\bar{\chi}(\mathop{\rm kish}N)$.
In the proof of Lemma 6.2.5 and below it, the double of a manifold $N$ with boundary is the manifold $\mathit{DN}$ obtained from $N\sqcup\overline{N}$, where $\overline{N}$ is a second copy of $N$, by identifying $\partial N\to\partial\overline{N}$ via the identity map. The following facts are standard and will be taken for granted. First, if $N$ is oriented then so is $\mathit{DN}$, by equipping $\overline{N}$ with the opposite orientation from that of $N$. Second, a hyperbolic structure with totally geodesic boundary on $N$ embeds isometrically into a boundaryless hyperbolic structure on $DN$ in which $\partial N$ hence sits as totally geodesic surface separating $N$ from $\overline{N}$.
Proof of Lemma 6.2.5.
Theorem 7.2 of [4] asserts that $\mathop{\rm vol}N\geq\frac{1}{2}V_{3}\|\mathit{DN}\|$, where $V_{3}$ is the volume of a regular ideal tetrahedron and $\|\cdot\|$ denotes the Gromov norm.
As used in the proof of [4, Th. 9.1], $\mathop{\rm kish}^{0}N$ admits a hyperbolic structure with totally geodesic boundary, which by Miyamoto’s universal lower bound on the volumes of manifolds with totally geodesic boundary [30, Th. 4.2] has volume at least $V_{\rm oct}\,\bar{\chi}(\mathop{\rm kish}N)$. Its double $D\mathop{\rm kish}^{0}N$ therefore has a finite-volume hyperbolic structure with volume at least twice this. And $D\mathop{\rm kish}^{0}N$ is the interior of the submanifold $D\mathop{\rm kish}N$ of $\mathit{DN}$ whose frontier is a disjoint union of incompressible tori. It thus follows that $V_{3}\|\mathit{DN}\|\geq\mathop{\rm vol}(D\mathop{\rm kish}^{0}N)$ from Theorem 6.5.5 of Thurston’s notes [35]. Combining these inequalities, we obtain
$\mathop{\rm vol}N\geq\frac{1}{2}V_{3}\|\mathit{DN}\|\geq\frac{1}{2}\mathop{\rm vol}(\textstyle{D\mathop{\rm kish}^{0}}N)\geq V_{\rm oct}\,\bar{\chi}(\mathop{\rm kish}N)$
Note that the present result’s conclusion asserts a strict inequality. If $\partial N$ is not totally geodesic then by [4, Th. 7.2] the leftmost inequality above is strict, and our desired conclusion holds. But if $\partial N$ is totally geodesic, then we claim that the rightmost inequality above is strict and again the desired conclusion holds; so in fact it holds unconditionally.
To prove the claim we note that since $\partial N$ is totally geodesic, the manifold $N$ is acylindrical. Hence, as pointed out in Subsection 6.1, we have $\mathop{\rm kish}N=N$; we therefore have $\mathop{\rm kish}^{0}N=N$. Theorem 4.2 of [30], applied to $N$, asserts for the rightmost inequality above that “equality holds only if $N$ is decomposed into $T^{n}(0)$’s”, where in this case (for $n=3$), the truncated regular simplex $T^{3}(0)$ of edgelength 0 is a regular ideal octahedron (compare [30, Example 5.1]). But any complete manifold that decomposes into copies of $T^{3}(0)$ has cusps, and the present $N$ is compact by hypothesis. Therefore the volume inequality is strict as claimed. ∎
The following result, Proposition 6.2.6, will be seen to be a direct consequence of deep results of Freedman–Hass–Scott [16] and Schoen–Yau [33]. This result, and its variant Proposition 6.2.7, will facilitate the applications of Lemma 6.2.5.
Proposition 6.2.6.
Let $S$ be a connected incompressible surface in a closed, orientable hyperbolic $3$-manifold $M$. The inclusion map $S\hookrightarrow M$ is homotopic to a least-area immersion $f\colon\thinspace S\to M$ that is either (i) an embedding, or (ii) a two-sheeted covering map to a one-sided surface $K$ isotopic to the core of a twisted $I$-bundle in $M$ bounded by $S$.
Proof.
Note that $M$, being hyperbolic, is $\mathbb{P}^{2}$-irreducible and aspherical. (These are topological hypotheses of the results of [33] and [16].) Since $S$ is incompressible, the inclusion map $S\hookrightarrow M$ is homotopic in $M$ to a smooth least-area immersion $f$, by the main result of [33]. Theorem 5.1 of [16] (which implicitly assumes connectedness of the domain $S$) then asserts that $f$ satisfies one of the alternatives (i), (ii). ∎
In Section 7 of [16], generalizations of the results of the kind that we have summarized in Proposition 6.2.6 are considered. The first four paragraphs of that section outline a proof that Proposition 6.2.6 remains true if the hypothesis that $M$ is closed is weakened. In particular, the proposition appears to remain true if $M$ is a compact hyperbolic $3$-manifold with smooth boundary, and the mean curvature of $\partial M$ with respect to the inward normal is everywhere non-negative. The arguments that are indicated depend on methods developed in [29]. Joel Hass has explained to us how these methods can be adapted to this purpose; the details appear to be rather involved. In this paper we need only the very special case in which $\partial M$ is connected and totally geodesic, and we prefer to provide a complete proof of this special result, stated as Proposition 6.2.7 below, that quotes only results that are proved in detail in the literature.
Proposition 6.2.7.
Let $S$ be a connected, incompressible (closed) surface in (the interior of) a compact, orientable hyperbolic $3$-manifold NN with connected totally geodesic boundary. Then one of the following alternatives holds:
  • (i-a)
    $S$ and $\partial N$ cobound a submanifold of $N$ which is a trivial $I$-bundle over a closed, connected surface;
  • (i-b)
    the inclusion map $S\hookrightarrow N$ is homotopic in $N$ to a least-area embedding $f\colon\thinspace S\to N$ whose image is contained in $\mathop{\rm int}N$; or
  • (ii)
    the inclusion map $S\hookrightarrow N$ is homotopic in $N$ to a least-area immersion which is a two-sheeted covering map to a one-sided surface $K\subset\mathop{\rm int}N$, isotopic to the core of a twisted $I$-bundle in $N$ bounded by $S$.
Proof of Proposition 6.2.7.
We first consider the case in which the inclusion map $S\hookrightarrow DN$ is homotopic in $DN$ to a map whose image is contained in $\partial N$. In this case, since $S$ and $\partial N$ are disjoint, it follows from [36, Lemma 5.3] that $S$ and $\partial N$ cobound a submanifold of $DN$ which is a trivial $I$-bundle over a closed, connected surface. Since $S$ is in the interior of $N$, this submanifold is contained in $N$. Thus Alternative (i-a) of the conclusion of the proposition holds in this case.
For the rest of the proof, we shall assume that:
6.2.7.1.
the inclusion map $S\hookrightarrow DN$ is not homotopic in $DN$ to a map whose image is contained in $\partial N$.
Proposition 6.2.6, applied with the double $\mathit{DN}\supset N$ playing the role of $M$, implies that the inclusion $S\hookrightarrow\mathit{DN}$ is homotopic in $\mathit{DN}$ to a least-area immersion $f\colon\thinspace S\to\mathit{DN}$ which is either (i) an embedding, or (ii) a two-sheeted covering map to a one-sided surface $K\subset DN$. (The second alternative of the conclusion of Proposition 6.2.6 is stronger than (ii), but the stronger information will not be used.)
The inclusion map $\partial N\hookrightarrow\mathit{DN}$ of the totally geodesic surface $\partial N$ is least-area, by Corollary 6.2.4, so since $S$ and $\partial N$ are disjoint, Theorem 6.2 of [16] further asserts that $f(S)$ is also disjoint from or identical to $\partial N$. In view of 6.2.7.1 we cannot have $f(S)=\partial N$. Hence:
(6.2.7.2) $f(S)\cap\partial N=\emptyset.$
Let us now fix a base point in $\partial N\subset N\subset DN$, and consider the covering space $p\colon\thinspace\widetilde{\mathit{DN}}\to\mathit{DN}$ determined by the image of the inclusion homomorphism $\pi_{1}(N)\to\pi_{1}(DN)$. There is a submanifold $N_{0}$ of $\widetilde{DN}$ which is mapped diffeomorphically onto $N$ by $p$; and since $\partial N$ is connected, the submanifold $Z\doteq\widetilde{DN}-\mathop{\rm int}N_{0}$ is also connected. We claim:
6.2.7.3.
The inclusion $\partial N_{0}\hookrightarrow Z$ is a homotopy equivalence.
[…]
In particular it follows from 6.2.7.3 that the inclusion $N_{0}\hookrightarrow\widetilde{DN}$ is a homotopy equivalence (which could be seen more directly).
Next, we claim:
6.2.7.4.
$f(S)\subset\mathrm{int}\,N$, and $f$ is homotopic in $\mathrm{int}\,N$ to the inclusion $S\hookrightarrow N$.
To prove 6.2.7.4, we first observe that the inclusion map $S\hookrightarrow N$ lifts to an embedding $j$ of $S$ in $\widetilde{DN}$. By the covering homotopy property of covering spaces, $j$ is homotopic in $\widetilde{DN}$ to some lift $\widetilde{f}$ of $f$. It follows from (6.2.7.2) that either $\widetilde{f}(S)\subset\mathop{\rm int}N_{0}$ or $\widetilde{f}(S)\subset\mathop{\rm int}Z$. But if $\widetilde{f}(S)$ were contained in $\mathop{\rm int}Z$, then by 6.2.7.3, $\widetilde{f}$ would be homotopic in $\widetilde{DN}$ to a map with image contained in $\partial N_{0}$, and hence $S\hookrightarrow\mathit{DN}$ would be homotopic in $\mathit{DN}$ to a map with image contained in $\partial N$; this contradicts 6.2.7.1. It follows that $\widetilde{f}(S)\subset\mathop{\rm int}N_{0}$. Now since $j$ and $\widetilde{f}$ are homotopic in $\widetilde{DN}$ and both map $S$ into $\mathop{\rm int}N_{0}$, and since we have observed that $N_{0}\hookrightarrow\widetilde{DN}$ is a homotopy equivalence, the maps $j$ and $\widetilde{f}$ are in fact homotopic in $\mathop{\rm int}N_{0}$. This immediately implies 6.2.7.4.
Since $f$ is a least-area map from $S$ to $DN$, and since $f(S)\subset N$ by 6.2.7.4, it follows from the definition that $f$ is a least-area map from $S$ to $N$. We also know from 6.2.7.4 that the least-area map $f:S\to N$ and the inclusion $S\hookrightarrow N$ are homotopic maps from $S$ to $N$.
If (i) holds, i.e. if $f$ is an embedding, then it follows from 6.2.7.4 that Alternative (i-b) of the present proposition holds.
Now suppose that (ii) holds. Thus $f$ is a two-sheeted covering map to a one-sided surface $K\subset DN$. According to 6.2.7.4 we have $K\subset\mathop{\rm int}N$. Let $J$ be a tubular neighborhood of $K$ in $\mathop{\rm int}N$. Then $J$ is a twisted $I$-bundle with core $K$. Now since $S$ is orientable and $K$ is not, the covering map $f:S\to K$ must be equivalent to the orientation covering of $K$. It follows that $f$ is homotopic in $J\subset\mathop{\rm int}N$ to a diffeomorphism $g:S\to\partial J$. But by 6.2.7.4, $f$ is homotopic in $\mathrm{int}\,N$ to the inclusion $S\hookrightarrow N$. Hence the maps $g$ and $S\hookrightarrow N$, which may both be regarded as embeddings of $S$ in $\mathop{\rm int}N$, are homotopic in $\mathop{\rm int}N$. It now follows from Corollary 5.5 of [36] that the surfaces $S$ and $\partial J=g(S)$ are isotopic in $N$.
Thus we may fix a self-diffeomorphism $h$ of $N$, isotopic to the identity, such that $h(S)=\partial J$. If we now set $J_{1}=h^{-1}(J)$ and $K_{1}=h^{-1}(K)$, then $J_{1}$ is a twisted $I$-bundle whose boundary is $S$, while $K_{1}$ is a core of $J_{1}$ and is isotopic to $K$. This gives Alternative (ii) of the conclusion of the present proposition. ∎
The next result is a variation on [Theorem 9.1] of [4] in which a strengthened hypotheses (that $M$ is compact rather than just having finite volume) yields the strengthened conclusion of a strict inequality.
Theorem 6.2.8.
[…]
(6.2.8.1) $\mathop{\rm vol}M>V_{\rm oct}\bar{\chi}(\mathop{\rm kish}(M\setminus\setminus S)).$
Proof.
The incompressibility of $S$, together with the hypotheses concerning the manifold $M$, implies that $S$ has genus at least $2$.
Since $S$ is incompressible, and $M$ has at most one boundary component, the hypotheses of either Proposition 6.2.6 or Proposition 6.2.7 hold. Hence either $\partial M=\emptyset$ and one of the alternatives (i), (ii) of the conclusion of Proposition 6.2.6 holds, or $\partial M$ is connected and one of the alternatives (i-a), (i-b) or (ii) of the conclusion of Proposition 6.2.7 holds.
First consider the case in which $\partial M$ is connected and Alternative (i-a) of Proposition 6.2.7 holds. Let $J$ denote the submanifold of $M$ which is a trivial $I$-bundle over a closed, connected surface and has boundary $S\cup\partial M$. Let $N$ denote the component of $M\setminus\setminus S$ that is distinct from $J$. Then since $\mathop{\rm kish}(J)=\emptyset$, and $N$ is diffeomorphic to $M$, we have
$\bar{\chi}(\mathop{\rm kish}(M\setminus\setminus S))=\bar{\chi}(\mathop{\rm kish}N)=\bar{\chi}(\mathop{\rm kish}M).$
Since $\partial M$ is by totally geodesic by hypothesis, it is least-area and hence minimal by Corollary 6.2.4. The conclusion thus follows from Lemma 6.2.5.
[…]
$\bar{\chi}(\mathop{\rm kish}N)=\bar{\chi}(\mathop{\rm kish}(M\setminus\setminus S)).$
The desired conclusion now follows from Lemma 6.2.5, upon noting that $\mathop{\rm vol}N=\mathop{\rm vol}M$. ∎
7. The capstone volume-topology dichotomy in the geodesic boundary case
In this section we prove this paper’s main result for hyperbolic $3$-manifolds with totally geodesic boundary, Theorem 7.2, which was stated in the Introduction. First, we strengthen Theorem 7.4 of [14] using Proposition 5.3 and Theorem 6.2.8.
Theorem 7.1.
Let $N$ be a compact, orientable hyperbolic $3$-manifold with $\partial N$ connected, totally geodesic, and of genus $2$, such that there is a $(1,1,1)$-hexagon in $\widetilde{N}$ and
$\cosh\ell_{1}<\frac{\cos(2\pi/9)}{2\cos(2\pi/9)-1}=1.43969$...
Then either the Heegaard genus $\mathrm{Hg}(N)$ is at most $4$ or $\mathrm{vol}(N)>2V_{\rm oct}$.
Recall from the Introduction that the quantity $V_{\rm oct}=3.6638$... referred to above is the volume of a regular ideal octahedron in $\mathbb{H}^{3}$, four times Catalan’s constant.
Proof.
We follow the proof of [14, Th. 7.4], replacing its first paragraph’s appeal to Propositions 6.8 and 6.9 of that paper by an appeal to Proposition 5.3 of the present paper in order to produce the non-degenerate trimonic submanifold $X$ of $N$ under the weaker bound on $\cosh\ell_{1}$. In the cases that $V=\overline{N-X}$ is boundary-reducible or of the form $|{\mathcal{W}}|$ for a book of $I$-bundles $\mathcal{W}$} (see 6.1), the previous proof’s topological arguments apply verbatim to show that $\mathrm{Hg}(N)\leq 4$.
The rest of the argument will be devoted to the case in which we have:
7.1.1.
$V\doteq\overline{N-X}$ is boundary-irreducible and is not of the form $|{\mathcal{W}}|$ for a book of $I$-bundles $\mathcal{W}$}.
In this case, we will slightly strengthen the geometric half of the conclusion’s dichotomy, from $\mathrm{vol}(N)>7.32$ in [14, Th. 7.4] to $\mathrm{vol}(N)>2V_{\rm oct}$ here. The proof of [14, Th. 7.4] appealed to Theorem 9.1 of [4]. To secure the improvement, we will appeal to Theorem 6.2.8 instead.
[…]
We conclude this section with the proof of the following result, which upgrades Theorem 1.1 of [14] and was stated in the Introduction.
Theorem 7.2.
Let $N$ be a compact, orientable hyperbolic $3$-manifold with $\partial N$ connected, totally geodesic, and of genus $2$. If $\mathrm{Hg}(N)\geq 5$ then $\mathrm{vol}(N)>2V_{\rm oct}$.
Proof.
For $N$ satisfying the Theorem’s hypotheses, if its universal cover $\widetilde{N}$ has no $(1,1,1)$-hexagon then by Corollary 4.7, $\mathrm{vol}(N)\geq 7.4>2V_{\rm oct}$. If $\widetilde{N}$ has a $(1,1,1)$-hexagon and the length $\ell_{1}$ of its shortest return path satisfies $\cosh\ell_{1}>1.439$, Theorem 4.5 gives the same lower bound on $\mathrm{vol}(N)$. We are left with the case that $\widetilde{N}$ does have a $(1,1,1)$-hexagon and $\cosh\ell_{1}\leq 1.439$. Here Theorem 7.1 implies that since $\mathrm{Hg}(N)\geq 5$, $\mathrm{vol}(N)>2V_{\rm oct}$.∎
8. Background for the results on closed manifolds
Definitions, Conventions, and Remarks 8.1.
The definitions, conventions and remarks given in Subsection 6.1 will be freely used in the rest of the paper, as will the ones given below.
We recall that the rank of a group $\Pi$ is the minimum cardinality of a generating set for $\Pi$.
As was mentioned in the Introduction, a group $\Pi$ is said to be $k$-free for a given positive integer $k$ if every subgroup of $\Pi$ whose rank is at most $k$ is free.
A group is said to be freely indecomposable if it is not trivial or infinite cyclic, and is not a free product of two non-trivial subgroups.
If $M$ is a compact, connected, orientable $3$-manifold, its Heegaard genus ${\rm Hg}(M)$ (see 6.1) is bounded below by $\mathop{{\rm rank}}\pi_{1}(M)$, which in turn is bounded below by $\dim H_{1}(M;{\mathbb{F}}_{p})$ for each prime $p$.
Proposition 8.2.
Let $M$ be a compact, orientable $3$-manifold-with-boundary, and let $F$ be a field. Then the dimension of the image of the inclusion homomorphism $H_{1}(\partial M;F)\to H_{1}(M;F)$ is equal to the sum of the genera of the components of $\partial M$.
Proof.
According to [21, Lemma 3.5], the dimension of the image of the boundary homomorphism $H_{2}(M,\partial M;{\mathbb{Q}})\to H_{1}(\partial M;{\mathbb{Q}})$ is one-half the dimension of $H_{1}(\partial M;{\mathbb{Q}})$. An examination of the proof of [Lemma 3.5] of [21] reveals that the proof goes through without change if $\mathbb{Q}$} is replaced by any field. Thus if $\partial:H_{2}(M,\partial M;F)\to H_{1}(\partial M;F)$ denotes the boundary homomorphism, the dimension of the image of $\partial$ is $(\dim H_{1}(\partial M;F))/2$. But we have $\dim H_{1}(\partial M;F)=2G$, where $G$ denotes the sum of the genera of the components of $\partial M$; furthermore, by the homology exact sequence of the pair $(M,\partial M)$, the image of $\partial$ is the kernel of the inclusion homomorphism $\iota:H_{1}(\partial M;F)\to H_{1}(M;F)$. Hence the kernel of $\iota$ has dimension $G$, and since the domain of $\iota$ is a vector space of dimension $2G$, the image of $\iota$ has dimension $G$ as well. ∎
Proposition 8.3.
Let $k\geq 3$ be an integer, and let $M$ be a closed, orientable, hyperbolic $3$-manifold. Suppose that $\dim H_{1}(M;{\mathbb{F}}_{2})\geq\max(3k-4,6)$, and that $\pi_{1}(M)$ is not $k$-free. Then $M$ contains a closed incompressible surface of some genus $g$ with $2\leq g\leq k-1$.
Proof.
Proposition 8.1 of [11] includes the fact that if $k\geq 3$ is an integer and $M$ is a closed, orientable hyperbolic $3$-manifold with $H_{1}(M;{\mathbb{F}}_{2})\geq\max(3k-4,6)$, then either $\pi_{1}(M)$ is $k$-free, or M contains a closed incompressible surface of genus at most k1k-1. Since the genus of an incompressible surface in $M$ is at least $2$ by 8.1, the present proposition follows. ∎
Proposition 8.4.
If $M$ is a closed, orientable, hyperbolic $3$-manifold and $\pi_{1}(M)$ is $3$-free, then $\mathop{\rm vol}M>3.08$.
Proof.
This is included in Corollary 9.3 of [3]. ∎
Lemma 8.5.
Let $M$ be a closed $3$-manifold. Set $V=H_{1}(M;{\mathbb{F}}_{2})$, and suppose that $P$ is a codimension-$2$ subspace of $V$. Set $k=dimP=(\dim V)-2$. Let $\widetilde{M}$ denote the regular covering of $M$, with covering group ${\mathbb{Z}}/2{\mathbb{Z}}\times{\mathbb{Z}}/2{\mathbb{Z}}$, that is determined by $P$ (so that the normal subgroup of $\pi_{1}(M)$ corresponding to $\widetilde{M}$ is the preimage of $P$ under the Hurewicz homomorphism $\pi_{1}(M)\to V$). Then $\dim H_{1}(\widetilde{M};{\mathbb{F}}_{2})\geq 2k+1$.
Proof.
Let $\Gamma_{1}$ denote the normal subgroup of $\pi_{1}(M)$ generated by all commutators and squares. Thus $\Gamma_{1}$ is the kernel of the Hurewicz homomorphism $\eta:\pi_{1}(M)\to H_{1}(M;{\mathbb{F}}_{2})$. According to the case $p=2$ of [34, Lemma 1.5], if $n$ is any integer less than or equal to $(\dim V)-2$, if $E$ is any subgroup generated by $n$ elements of $\pi_{1}(M)$, and if $D$ denotes the subgroup $D\doteq E\Gamma_{1}$ of $\pi_{1}(M)$, we have $\dim H_{1}(D;{\mathbb{F}}_{2})\geq 2n+1$. To apply this, we take $n=k=(\dim V)-2$, we choose elements $x_{1},\ldots,x_{k}$ of $\pi_{1}(M)$ whose images under $\eta$ form a basis of $P$, and we take $E$ to be the subgroup of $\pi_{1}(M)$ generated by $x_{1},\ldots,x_{k}$. Then $D\doteq E\Gamma_{1}=\eta^{-1}(P)$, and hence $H_{1}(\widetilde{M};{\mathbb{F}}_{2})$ is isomorphic to $H_{1}(D;{\mathbb{F}}_{2})$. The result now follows. ∎
9. Incompressible surfaces, homology rank, and volume
Lemma 9.1.
If $\mathcal{W}$} is a connected book of $I$-bundles (see 6.1), each of whose pages has negative Euler characteristic, then $\dim H_{1}(|{\mathcal{W}}|)\leq 2\bar{\chi}(|{\mathcal{W}}|)+1$.
Proof.
This is Lemma 2.11 of [3]. (The connectedness hypothesis is missing from the statement of [3, Lemma 2.11], but it is used in the proof, and holds in the context of the applications given in [3]. ∎
Lemma 9.2.
Let $g$ be a positive integer, and $M$ be a closed, orientable hyperbolic $3$-manifold that contains a closed, connected incompressible surface of genus $g$. Suppose that ${\rm Hg}(M)>2g+1$. Then there exist a closed, connected incompressible surface $S\subset M$ such that either
  1. (1)
    $\bar{\chi}(\mathop{\rm kish}(M\setminus\setminus S))\geq 2$, or
  2. (2)
    the surface $S$ separates $M$, and $M\setminus\setminus S$ has an acylindrical component.
Proof.
According to [18, Proposition 13.2], the hypotheses imply that there is a connected incompressible surface $S\subset M$ such that either Alternative (2) of the statement of the present lemma holds, or one of the following alternatives holds:
  • (1a)
    the surface $S$ separates $M$, and for each component $B$ of $M\setminus\setminus S$ we have $\mathop{\rm kish}(B)\neq\emptyset$; or
  • (1b)
    the surface $S$ does not separate $M$, and $\bar{\chi}(\mathop{\rm kish}(M\setminus\setminus S))\geq 2h-2$, where $h$ denotes the genus of $S$.
(Proposition 13.2 of [18] also gives information about the genus of $S$, which will not be needed here.)
If Alternative (1a) holds then $\mathop{\rm kish}(M\setminus\setminus S)$ has at least two components, and according to 6.1 we have $\bar{\chi}(K)\geq 1$ for each component $K$ of $\mathop{\rm kish}(M\setminus\setminus S)$. Hence Alternative (1) of the conclusion of the present lemma holds in this case. If Alternative (1b) holds, then since the genus $h$ of $S$ is at least $2$ by 8.1, we have $\bar{\chi}(\mathop{\rm kish}(M\setminus\setminus S)\geq 2$, i.e. Alternative (1) of the present lemma holds in this case as well. ∎
The following result, like Theorem 6.2.8, is proved using the techniques of [4].
Proposition 9.3.
Let $S$ be a connected, incompressible surface in a closed, orientable hyperbolic $3$-manifold $M$ and let $A$ be an acylindrical component of $M\setminus\setminus S$. Then $A$ is diffeomorphic to a hyperbolic manifold $N$ with totally geodesic boundary, and $\mathop{\rm vol}M\geq\mathop{\rm vol}N$.
Proof.
This is a formal consequence of Propositions 6.1 and 6.2 of [8]. Note that if $M$, $S$ and $A$ satisfy the hypotheses of Proposition 9.3, then $A$ is irreducible, boundary-irreducible and acylindrical. In Section 6 of [8], a certain real-valued invariant $\mathop{\rm geodvol}A$ is defined for a compact, connected $3$-manifold AA with non-empty boundary. Proposition 6.1 of [8] asserts that if $A$ is irreducible, boundary-irreducible and acylindrical, then $A$ is diffeomorphic to a hyperbolic manifold $N$ with totally geodesic boundary, and $\mathop{\rm vol}N=\mathop{\rm geodvol}A$. Proposition 6.2 of [8] asserts that if $A$ is an acylindrical component of $M\setminus\setminus S$, where $S$ is a connected, incompressible surface in a closed, orientable hyperbolic $3$-manifold $M$, then $\mathop{\rm vol}M\geq\mathop{\rm geodvol}A$. Proposition 9.3 now follows immediately.
(The actual definition of $\mathop{\rm geodvol}A$ is that it is one-half the Gromov volume of the double of $A$. This definition of course enters into the proofs of Propositions 6.1 and 6.2 of [8].) ∎
Proposition 9.4.
Let $M$ be a closed, orientable, hyperbolic $3$-manifold, let $k\geq 3$ be an integer, and suppose that $\dim H_{1}(M;{\mathbb{F}}_{2})\geq\max(3k-4,6)$. Then either $\pi_{1}(M)$ is $k$-free, or $\mathop{\rm vol}M>2V_{\rm oct}$.
Proof.
Set $r=\dim H_{1}(M;{\mathbb{F}}_{2})$.
We shall assume that $\pi_{1}(M)$ is not $k$-free, and show that $\mathop{\rm vol}M>2V_{\rm oct}$. Since $\pi_{1}(M)$ is not $k$-free and $r\geq\max(3k-4,6)$, it follows from Proposition 8.3 that $M$ contains a closed incompressible surface of some genus $g$ with $2\leq g\leq k-1$.
Since $k\geq g+1$, we have $r\geq\max(3k-4,6)>2g+1$. In particular, we have ${\rm Hg}(M)>2g+1$ (see 8.1). Hence by Lemma 9.2, there is a closed, connected incompressible surface $S\subset M$ such that either
  1. (1)
    $\bar{\chi}(\mathop{\rm kish}(M\setminus\setminus S))\geq 2$, or
  2. (2)
    the surface $S$ separates $M$, and $M\setminus\setminus S$ has an acylindrical component.
If (1) holds, Theorem 6.2.8 gives $\mathop{\rm vol}M>2V_{\rm oct}$, so that the conclusion of the lemma is true in this case. For the rest of the proof, we shall assume that (2) holds but that (1) does not.
Fix an acylindrical component $A$ of $M\setminus\setminus S$, and let $B$ denote the other component of $M\setminus\setminus S$. Since $A$ is acylindrical, it follows from Proposition 9.3 that $A$ is diffeomorphic to a hyperbolic $3$-manifold NN with totally geodesic boundary, and that
(9.4.1) $\mathop{\rm vol}M\geq\mathop{\rm vol}N.$
According to 6.1, we have $\bar{\chi}(K)\geq 1$ for each component $K$ of $\mathop{\rm kish}(M\setminus\setminus S)$; thus $\bar{\chi}(\mathop{\rm kish}(M\setminus\setminus S))$ is bounded below by the number of components of $\mathop{\rm kish}(M\setminus\setminus S)$. Since $A$ is acylindrical, we have $\mathop{\rm kish}(A)=A\neq\emptyset$ by 6.1. If $\mathop{\rm kish}(B)$ were also non-empty, it would follow that $\mathop{\rm kish}(M\setminus\setminus S)$ had at least two components, and therefore that $\bar{\chi}(\mathop{\rm kish}(M\setminus\setminus S))\geq 2$; this would mean that Alternative (1) above holds, a contradiction. Hence $\mathop{\rm kish}(B)=\emptyset$.
Consider the subcase in which $S$ has genus at least $3$. If $h$ denotes the genus of $S$, it follows from Theorem 5.4 of [30] that $\mathop{\rm vol}N$ is bounded below by $h\mathop{\rm vol}T_{\pi/(3h)}$, where $T_{\theta}$ denotes a truncated regular simplex of dihedral angle $\theta$, in the sense defined in [30]. Proposition 1.1 of [30] gives a formula for the volume of $T_{\theta}$ which is visibly monotone decreasing in $\theta$. Since $h\geq 3$ it follows that $\mathop{\rm vol}N\geq h\mathop{\rm vol}T_{\pi/(3h)}\geq 3\mathop{\rm vol}T_{\pi/9}=10.4\ldots>2V_{\rm oct}$, and the lemma is established in this subcase.
There remains the subcase in which $S$ has genus $2$. In the following argument, all homology groups in this argument will be understood to have coefficients in ${\mathbb{F}}_{p}$.
Since $\mathop{\rm kish}(B)=\emptyset$, it follows from an observation made in 6.1 that the connected manifold $B$ is the underlying manifold of a book of $I$-bundles $\mathcal{W}$}, each of whose pages has negative Euler characteristic. Hence Lemma 9.1 gives $\dim H_{1}(B)\leq 1+2\bar{\chi}(B)$. But since $S$ has genus $2$ we have $\bar{\chi}(S)=2$, and since $S=\partial B$ we have $\bar{\chi}(B)=\bar{\chi}(S)/2=1$. Hence $\dim H_{1}(B)\leq 3$.
Consider the Mayer-Vietoris fragment
${H_{1}(S)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{j}$${H_{1}(A)\oplus H_{1}(B)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H_{1}(M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H_{0}(S)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\tau}$${H_{0}(A)\oplus H_{0}(B).}$
The homomorphism $\tau$ is injective since $S$ is connected and $A\neq\emptyset$. Hence if $J$ denotes the image of $j$, we have an exact sequence
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${J\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{j}$${H_{1}(A)\oplus H_{1}(B)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H_{1}(M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
The exactness of the latter sequence implies that
(9.4.2) $\dim H_{1}(A)=\dim J+\dim H_{1}(M)-\dim H_{1}(B).$
We have seen that $\dim H_{1}(B)\leq 3$, and by definition we have $\dim H_{1}(M)=r$. To estimate $\dim J$, we recall that the homomorphism $j:H_{1}(S)\to H_{1}(A)\oplus H_{1}(B)$ is defined by $j(x)=(\iota_{A}(x),\iota_{B}(x))$, where $\iota_{A}:H_{1}(S)\to H_{1}(A)$ and $\iota_{B}:H_{1}(S)\to H_{1}(B)$ are the inclusion homomorphisms. Hence $\dim J$ is bounded below by the dimension of the image of $\iota_{A}$. According to Proposition 8.2, the dimension of the image of $\iota_{A}$ is equal to the genus of $S$, which is $2$. Thus $\dim J\geq 2$, and (9.4.2) implies that $\dim H_{1}(A)\geq 2+r-3=r-1$, or equivalently that $\dim H_{1}(N)\geq r-1$.
In particular, we have ${\rm Hg}(N)\geq r-1$ (see 8.1). Since $r\geq 6$ by hypothesis, ${\rm Hg}(N)$ is in particular at least $5$. According to Theorem 7.2, this implies that $\mathop{\rm vol}N>2V_{\rm oct}$. Since $\mathop{\rm vol}M\geq\mathop{\rm vol}N$ by (9.4.1), the conclusion of the lemma follows in this final subcase. ∎
10. Homology of manifolds with volume at most $V_{\rm oct}/2$
The following result was stated in the introduction.
Theorem 10.1.
Let $M$ be a closed, orientable hyperbolic $3$-manifold with $\mathop{\rm vol}M\leq V_{\rm oct}/2$. Then $\dim H_{1}(M;{\mathbb{F}}_{2})\leq 4$.
Proof.
We set $\Pi=\pi_{1}(M)$, and $V=H_{1}(M;{\mathbb{F}}_{2})$. We identify $V$ with $H_{1}(\Pi;{\mathbb{F}}_{2})$.
We shall assume that $\dim V\geq 5$ and show that $\mathop{\rm vol}M>V_{\rm oct}/2$, thus proving the theorem.
[…]
11. Volumes of manifolds with small cup product rank
The next two lemmas are needed for the proof of Theorem 11.3, which was stated in the introduction.
Lemma 11.1.
Suppose that $M$ is a closed, aspherical 33–manifold. Set $r=\dim H_{1}(M;{\mathbb{F}}_{2})$, and let $t$ denote the dimension of the image of the cup product pairing $H^{1}(M;{\mathbb{F}}_{2})\otimes H^{1}(M;{\mathbb{F}}_{2})\to H^{2}(M;{\mathbb{F}}_{2})$. Then for any two-sheeted covering $\widetilde{M}$ of $M$, we have $\dim H_{1}(M;{\mathbb{F}}_{2})\geq 2r-t-1$.
Proof.
This is the case $m=1$ of [10, Proposition 3.5]. ∎
Lemma 11.2.
Let $M$ be a closed, orientable hyperbolic $3$-manifold. Suppose that $\pi_{1}(M)$ is $4$-free and that $\dim H_{1}(M;{\mathbb{F}}_{2})\geq 6$. Then $\mathop{\rm vol}(M)>3.69$.
Proof.
This follows from the proof of [18, Proposition 14.5]. The latter proposition is equivalent to the statement that if $M$ is a closed, orientable hyperbolic $3$-manifold such that $\dim H_{1}(M;{\mathbb{F}}_{2})\geq 8$, then $\mathop{\rm vol}(M)>3.69$. In the first two paragraphs of the proof of [18, Proposition 14.5], it is shown that the hypothesis $\dim H_{1}(M;{\mathbb{F}}_{2})\geq 8$ implies that either $M$ contains an incompressible surface of genus at most $3$, or $\pi_{1}(M)$ is $4$-free, and it is shown that in the former case one has a stronger conclusion than $\mathop{\rm vol}(M)>3.69$. The remainder of the proof is devoted to the case in which $M$ satisfies the homological hypothesis and $\pi_{1}(M)$ is $4$-free. However, an examination of this part of the proof reveals that while the assumption of $4$-freeness is used in an essential way, the only homological information that is used is that $\dim H_{1}(M;{\mathbb{F}}_{2})\geq 6$. Thus this portion of the proof establishes the present lemma.
More specifically, the argument given in [18, Proposition 14.5] for the $4$-free case is divided into four subcases, labeled (a)–(d). The proofs of the inequality $\mathop{\rm vol}(M)>3.69$ in Subcases (a) and (d) make strong use of $4$-freeness, but do not depend on any homological information. The proofs in Subcases (b) and (c) are direct applications of Lemma 14.4 of [18]. That lemma does not involve $4$-freeness, but does have a homological hypothesis, namely $\dim H_{1}(M;{\mathbb{F}}_{2})\geq 6$. ∎
As we mentioned in the Introduction, the following theorem improves on Theorem 1.2 of [14]; and the proof given here, besides strengthening the result, provides more details than the proof in [14] and corrects a citation.
Theorem 11.3.
Let $M$ be a closed, orientable hyperbolic $3$-manifold. Set $r=\dim H_{1}(M;{\mathbb{F}}_{2})$, and let $t$ denote the dimension of the image of the cup product pairing $H^{1}(M;{\mathbb{F}}_{2})\otimes H^{1}(M;{\mathbb{F}}_{2})\to H^{2}(M;{\mathbb{F}}_{2})$. Then:
  1. (1)
    if $r\geq 5$ and $t\leq 1$, we have $\mathop{\rm vol}(M)>3.57$; and
  2. (2)
    if $r\geq 6$ and $t\leq 3$, or $r\geq 7$ and $t\leq 5$, we have $\mathop{\rm vol}(M)>V_{\rm oct}$.
Proof.
We shall first prove Assertions (1) and (2) in the case where $\pi_{1}(M)$ is $4$-free. In this case Assertion (1) is an immediate consequence of Proposition 12.12 of [18], which asserts that every closed, orientable hyperbolic $3$-manifold with $4$-free fundamental group has volume greater than $3.57$. To prove Assertion (2) in this case, we note that the hypothesis of Assertion (2) implies that $r\geq 6$, which by Lemma 11.2 above and the $4$-freeness of $\pi_{1}(M)$ implies $\mathop{\rm vol}(M)>3.69>V_{\rm oct}$.
[…]